### v1

# On quasi-Albanese morphisms for log canonical Calabi-Yau pairs

Abstract.
We study the quasi-Albanese morphisms for log canonical Calabi-Yau pairs.
2020 Mathematics Subject Classification:
14E30
1. Introduction
Let $X$ be a complex projective variety with canonical singularities and $K_{X}$ torsion. Kawamata proved that its Albanese morphism $\alpha:X\to A$ is an étale fiber bundle [Kaw85, Theorem 8.3]. Inspired by the works [Iit76, Kaw81, Zha05, Fuj24, BFPT24], we study quasi-Albanese morphisms for (the open part of) log canonical Calabi-Yau pairs.
Before stating the main theorem, we recall the definition of logarithmic Kodaira dimension and a theorem by Kawamata and Fujino. Let $X^{0}$ be a smooth quasi-projective variety. Let $X$ be a compactification of $X^{0}$ such that $D_{X}:=X\setminus X^{0}$ is a simple normal crossing divisor. The logarithmic Kodaira dimension $\overline{\kappa}(X^{0})$ of $X^{0}$ is defined as $\kappa(K_{X}+D_{X})$. By Iitaka [Iit77], $\overline{\kappa}(X^{0})$ is independent of the choice of such $X$.
Theorem 1.
[Kaw81, Thoerem 28], see also [Fuj24, Theorem 1.2] If $\overline{\kappa}(X^{0})=0$, then the quasi-Albanese morphism of $X^{0}$ is dominant and has irreducible general fibers.
The following is our main theorem.
Theorem 2.
Let $(X,D_{X})$ be a log smooth pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$. Let $X^{0}=X\setminus D_{X}$ and let $\alpha:X^{0}\to G$ be its quasi-Albanese morphism. Then
  1. (1)
    $\alpha$ is surjective in codimension one: for any prime divisor $P\subset G$, there is a prime divisor $Q\subset X^{0}$ dominating $P$,
  2. (2)
    $\alpha$ is flat in codimension one: for any prime divisor $Q\subset X^{0}$, if $Q$ do not dominate $G$, then $\alpha(Q)$ is of codimension one,
  3. (3)
    $\alpha$ is semistable in codimension one: for any prime divisor $P\subset G$, the pair $(X,\alpha^{*}P)$ is log canonical near the generic point of $P$.
Theorem 2 is a consequence of the following theorem, whose proof depends on Theorem 1 and the techniques from the minimal model program.
Theorem 3.
Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$. Set $q(X):=\dim H^{1}(O_{X})$ and $q(X,D_{X}):=\dim H^{0}(\Omega_{X^{0}}^{1}(\log D_{X}|_{X^{0}}))$, where $X^{0}$ is the log smooth locus of $(X,D_{X})$. Let $A$ be the Albanese variety of $X$ and let $G$ be the quasi-Albanese variety of $X\setminus D_{X}$. Then there are
  • a $\mathbb{Q}$-factorial dlt pair $(W,D_{W})$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$,
  • a crepant birational contraction $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ with $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$,
  • $(q(X,D_{X})-q(X))$’s line bundles $L_{i}\in\mathrm{Pic}^{0}(A)$, and
  • a commutative diagram,
[…]
    where $\mathbb{P}_{A}:=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1})$ is a compactification of $G$, $H=\mathbb{P}_{A}\setminus G$ is a simple normal crossing divisor, $g$ is a fibration, and $p$ is the natural projection,
such that the following holds.
  1. (1)
    $p\circ g$ is the Albanese morphism of $W$, $h$ is the quasi-Albanese morphism of $W\setminus D_{W}$.
  2. (2)
    The $g$-vertical part of $D_{W}$ coincides with $g^{-1}H$.
  3. (3)
    The canonical bundle formula for $g:(W,D_{W})\to\mathbb{P}_{A}$ has the form
    $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H),$
    and the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable.
As a corollary, if $q(X)=0$ and $q(X,D_{X})=\dim X$, then $(X,D_{X})$ is crepant birational to $(\mathbb{P}^{n},(\prod_{i}z_{i}=0))$. Another corollary is that if $q(X,D_{X})-q(X)\geq 2$, then $D_{X}$ is connected.
In the previous Theorem, the pair $(W,D_{W})$ is closely related to $(X,D_{X})$. For example, they have the same orbifold fundamental group, and the same dual complex (up to a PL-homeomorphism) (cf. [dFKX17, Proposition 11]). On the other hand, the fibration $g:(W,D_{W})\to\mathbb{P}_{A}$ has nice properties, which may reduce some questions about $(X,D_{X})$ to the case $q(X,D_{X})=0$ by induction on dimension. We give an example in the last section.
Acknowledgments
The author thanks professors Zhan Li, Lei Zhang, and Junchao Shentu for useful discussions. The author is supported by USTC.
2. Preparations
We work over the field of complex numbers. For the singularities of pairs and the minimal model program, we refer to the book [KM98].
A projective surjective morphism between normal varieties with connected fibers is called a fibration. Let $f:X\to Y$ be a fibration between normal varieties. Let $D$ be an effective $\mathbb{Q}$-divisor on $X$. We can write $D=D^{h}+D^{v}$, where $D^{v}$ does not dominate $Y$, and each component of $D^{h}$ dominates $Y$. $D^{v}$ (resp. $D^{h}$) is called the vertical part (resp. horizontal) part of $D$. $D$ is called vertical (resp. horizontal) if $D=D^{v}$ (resp. $D=D^{h}$). Assume that $D$ is vertical, then $D$ is called exceptional over $Y$ if $\operatorname{codim}f(D)\geq 2$, and very exceptional over $Y$ if for any prime divisor $P\subset Y$, there is a prime divisor $Q\subset X$ such that $Q$ dominates $P$ but $Q\not\subset D$.
For a birational map $\pi:X\dashrightarrow Y$ between normal varieties, we set $\operatorname{Ex}(\pi):=\sum P$, where the sum is taken over all prime divisors on $X$ which are exceptional over $Y$. A birational map $\pi:X\dashrightarrow Y$ is called a birational contraction if $\operatorname{Ex}(\pi^{-1})=0$.
We will use the following fact implicitly: Let $A$, $B$ be two effective $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisors on a projective variety $X$. If $\operatorname{Supp}A=\operatorname{Supp}B$, then $\kappa(A)=\kappa(B)$. If $\operatorname{Supp}A\subset\operatorname{Supp}B$, then $\kappa(A)\leq\kappa(B)$. Let $f:Y\to X$ be a projective surjective morphism from another projective variety, then $f^{-1}\operatorname{Supp}A=\operatorname{Supp}f^{*}A$.
For a projective variety $X$, we set $q(X):=\dim H^{1}(O_{X})$, and for a projective lc pair $(X,D_{X})$, we set $q(X,D_{X}):=\dim H^{0}(\Omega_{X^{0}}^{1}(\log\lfloor D_{X}|_{X^{0}}\rfloor))$, where $X^{0}$ is the log smooth locus of $(X,D_{X})$.
2.1. Differential forms on lc pairs
Theorem 4 ([GKKP11]).
Let $(X,D_{X})$ be an lc pair. Let $X^{0}$ be its log smooth locus and let $\mathrm{nonklt}(X,D_{X})$ be its non-klt locus. Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}$ be the codimension one part of $\pi^{-1}\mathrm{nonklt}(X,D_{X})$. Then $\pi_{*}\Omega_{Y}^{k}(\log D_{Y})$ is reflexive and $H^{0}(Y,\Omega_{Y}^{k}(\log D_{Y}))\cong H^{0}(X^{0},\Omega_{X^{0}}^{k}(\log\lfloor D_{X}|_{X^{0}}\rfloor))$ for all $k$.
Proposition 5.
Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced. Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Then
  1. (1)
    $q(X)=q(Y)$,
  2. (2)
    $q(X,D_{X})=q(Y,D_{Y})$,
  3. (3)
    $\kappa(K_{X}+D)=\kappa(K_{Y}+D_{Y})$ and $\nu(K_{X}+D)=\nu(K_{Y}+D_{Y})$,
where $\kappa$ and $\nu$ are Kodaira dimension and numerical Kodaira dimension respectively.
Proof.
Since klt varieties have rational singularities, we have (1). Since $(X,D_{X})$ is lc, there is an effective exceptional divisor $E$ on $Y$ such that $K_{Y}+D_{Y}=\pi^{*}(K_{X}+D_{X})+E$, so we have (3).
Now we prove (2). Note that $D_{Y}\geq\pi^{-1}D_{X}$ and $D_{X}=\mathrm{nonklt}(X,D_{X})$. By Theorem 4, to show (2), it suffices to show that the injective homomorphism
$\iota:H^{0}(\Omega_{Y}^{1}(\log\pi^{-1}D_{X}))\to H^{0}(\Omega_{Y}^{1}(\log D_{Y}))$
is an isomorphism. Write $\pi^{-1}D_{X}=\sum_{i}P_{i}$ and $D_{Y}-\pi^{-1}D_{X}=\sum_{j}E_{j}$ as sum of prime divisors. By the exact sequence
$0\to\Omega_{Y}^{1}\to\Omega_{Y}^{1}(\log(\sum_{i}P_{i}+\sum_{j}E_{j}))\to\oplus_{i}O_{P_{j}}\oplus\oplus_{j}O_{E_{j}}\to 0,$
one checks that $\iota$ is an isomorphism iff $\sum_{i}a_{i}c_{1}(O_{Y}(P_{i}))+\sum_{j}b_{j}c_{1}(O_{Y}(E_{j}))=0$ implies that all $b_{j}$’s are zero. Which is indeed the case since all $E_{j}$’s are exceptional and $X$ is $\mathbb{Q}$-factorial.∎
Lemma 6.
Let $(X,D_{X})$ and $(Y,D_{Y})$ be two projective $\mathbb{Q}$-factorial dlt pairs with reduced $D_{X}$ and $D_{Y}$. Assume that there is a birational contraction $\pi:X\dashrightarrow Y$ such that $D_{X}=\pi^{-1}_{*}D_{Y}+\operatorname{Ex}(\pi)$ then $q(X,D_{X})=q(Y,D_{Y})$.
Proof.
Let $X\xleftarrow{p}W\xrightarrow{q}Y$ be a common log resolution. Then $q^{-1}_{*}D_{Y}+\operatorname{Ex}(q)=p^{-1}_{*}D_{X}+\operatorname{Ex}(p)$. Hence, the Lemma follows from the previous proposition. ∎
2.2. Quasi-Albanese morphism
Theorem 7 (Quasi-Albanese morphism [Iit76], see also [Fuj24]).
Let $X$ be a smooth quasi-projective variety. Then there is a morphism $\alpha:X\to G$ to a quasi-abelian variety such that
(1) for any morphism $\alpha^{\prime}:X\to G^{\prime}$ to a quasi-abelian variety, there is a morphism $h:G\to G^{\prime}$ such that $\alpha^{\prime}=h\circ\alpha$, and
(2) $\alpha$ is uniquely determined.
8Construction.
We recall the constructions of quasi-Albanese morphisms in [Iit76, Fuj24]. Let $(X,D_{X}=\sum_{i=1}^{m}D_{i})$ be a projective log smooth pair with $D_{X}$ reduced. Where $D_{i}$’s are prime divisors. Let $X^{0}:=X\setminus D_{X}$ and let $\iota:X^{0}\to X$ be the open embedding. Let $p:X\to A$ be the Albanese morphism. Set $q=q(X)$, $\overline{q}=q(X,D_{X})$, and $d=\overline{q}-q$.
Let $\eta_{i}$, $1\leq i\leq d$, be a basis of the free part of $\ker(\iota_{*}:H_{1}(X^{0},\mathbb{Z})\to H_{1}(X,\mathbb{Z}))$, and let $\xi_{i}\in H_{1}(X^{0},\mathbb{Z})$, $1\leq i\leq 2q$, be such that these $\eta_{i}$’s and $\xi_{i}$’s form a basis of the free part of $H_{1}(X^{0},\mathbb{Z})$.
Let $\omega_{i}$, $1\leq i\leq q$, be a basis of $H^{0}(\Omega_{X}^{1})$, and let $\phi_{i}\in H^{0}(\Omega_{X}^{1}(\log D))$, $1\leq i\leq d$, be such that these $\omega_{i}$’s and $\phi_{i}$’s form a basis of $H^{0}(\Omega_{X}^{1}(\log D_{X}))$, and satisfy $\int_{\eta_{i}}\omega_{j}=0$ and $\int_{\eta_{i}}\phi_{j}=\delta_{ij}$ (the Kronecker delta).
Set
$A_{i}:=(\int_{\xi_{i}}\omega_{1},\cdots,\int_{\xi_{i}}\omega_{q},\int_{\xi_{i}}\phi_{1},\cdots,\int_{\xi_{i}}\phi_{d})=:(\hat{A}_{i},C_{i}),1\leq i\leq 2q,\text{ and }$
$B_{i}:=(\int_{\eta_{i}}\omega_{1},\cdots,\int_{\eta_{i}}\omega_{q},\int_{\eta_{i}}\phi_{1},\cdots,\int_{\eta_{i}}\phi_{d})=:(0,\hat{B}_{i}),1\leq i\leq d.$
The quasi-Albanese variety of $X^{0}$ is given by the quasi-abelian variety
$G:=\mathbb{C}^{\overline{q}}/\sum_{i}\mathbb{Z}A_{i}+\sum_{j}\mathbb{Z}B_{j}.$
We have the following Chevalley decomposition of $G$:
$0\to\mathbb{C}^{d}/\sum_{j}\mathbb{Z}\hat{B}_{j}\cong(\mathbb{C}^{*})^{d}\to G\to\mathbb{C}^{q}/\sum_{i}\mathbb{Z}\hat{A}_{i}=A\to 0,$
and $A$ coincides with the Albanese variety of $X$. The isomorphism $\mathbb{C}^{d}/\sum_{j}\mathbb{Z}\hat{B}_{j}\cong(\mathbb{C}^{*})^{d}$ is given by
$(x_{1},\cdots,x_{d})\mapsto(\exp(2\pi\mathrm{i}x_{1}),\cdots,(\exp(2\pi\mathrm{i}x_{d})).$
Fix a point $x_{0}\in X^{0}$. The quasi-Albanese morphism of $X^{0}$ is then given by
$X^{0}\to G,x\mapsto\left(\int_{x_{0}}^{x}\omega_{1},\cdots,\int_{x_{0}}^{x}\omega_{q},\int_{x_{0}}^{x}\phi_{1},\cdots,\int_{x_{0}}^{x}\phi_{d}\right).$
Remark 9.
Keep the notations introduced in the previous construction. Let $G^{\prime}\to\mathbb{C}^{q}$ be the pullback of the $(\mathbb{C}^{*})^{d}$-bundle $G\to A$ by the cover $\mathbb{C}^{q}\to A$. Then we have $G^{\prime}=\mathbb{C}^{\overline{q}}/\sum_{i}\mathbb{Z}B_{i}\cong\mathbb{C}^{q}\times(\mathbb{C}^{*})^{d}$ and
$G=\mathbb{C}^{q}\times(\mathbb{C}^{*})^{d}/\pi_{1}(A),$
where $\pi_{1}(A)=\sum_{i}\mathbb{Z}{\hat{A}_{i}}$ acts on $\mathbb{C}^{q}\times(\mathbb{C}^{*})^{d}$ diagonally by the group homomorphism:
$\pi_{1}(A)\to\operatorname{Aut}((\mathbb{C}^{*})^{d}),\hat{A}_{i}\mapsto\left((t_{1},\cdots,t_{d})\mapsto(t_{1}\exp(2\pi\mathrm{i}\int_{\xi_{i}}\phi_{1}),\cdots,t_{d}\exp(2\pi\mathrm{i}\int_{\xi_{i}}\phi_{d}))\right),1\leq i\leq q.$
For each $j$, let $L_{j}\in\mathrm{Pic}^{0}(A)$ be the line bundle corresponding to the representation
(1) $\pi_{1}(A)\to\mathbb{C}^{*},\hat{A}_{i}\mapsto\exp(2\pi\mathrm{i}\int_{\xi_{i}}\phi_{j}),1\leq i\leq q.$
Then $G$ is the fiber product over $A$ of the total spaces of $L_{j}$’s with zero sections removed, in other words,
$G=\operatorname{Spec}_{A}\bigoplus_{(m_{1},\cdots,m_{d})\in\mathbb{Z}^{d}}L_{1}^{-m_{1}}\otimes\cdots\otimes L_{d}^{-m_{d}}.$
$G$ has the following natural compactification:
(2) $G\subset\mathbb{P}_{A}:=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1}).$
Let $a_{ij}=2\pi\mathrm{i}\mathrm{Res}_{D_{i}}\phi_{j}$, $1\leq i\leq m$, $1\leq j\leq d$. For each $j$, we have the following.
  1. (1)
    $a_{ij}\in\mathbb{Z}$ for all $i$.
  2. (2)
    $p^{*}L_{j}$ is given by the representation $\pi_{1}(X)\to\mathbb{C}^{*},\gamma\mapsto\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})$.
  3. (3)
    $t_{j}(x):=\exp(2\pi\mathrm{i}\int_{x_{0}}^{x}\phi_{j})$ is a rational section of $p^{*}L_{j}$ with $\mathrm{div}(t_{j})=\sum_{i}a_{ij}D_{i}$. In particular, $p^{*}L_{j}\cong O_{X}(\sum_{i}a_{ij}D_{i})$.
Proof.
For each $i$, let $\delta_{i}$ be a small loop around $D_{i}$.
(1) Since $\delta_{i}\in\ker(H_{1}(X^{0},\mathbb{Z})\to H_{1}(X,\mathbb{Z}))$, we have $\delta_{i}=\sum_{k}n_{k}\eta_{k}+t$ for some integers $n_{k}$ and a torsion element $t$. By the residue theorem, we have
$2\pi\mathrm{i}\mathrm{Res}_{D_{i}}\phi_{j}=\int_{\delta_{i}}\phi_{j}=\sum_{k}n_{k}\int_{\eta_{k}}\phi_{j}+\int_{t}\phi_{j}=\sum_{k}n_{k}\delta_{kj}=n_{j}\in\mathbb{Z}.$
(2) By the expression $\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})$, we mean $\exp(2\pi\mathrm{i}\int_{\gamma^{\prime}}\phi_{j})$, where $\gamma^{\prime}$ is any other loop in $X^{0}$ which is homotopic to $\gamma$. Since $\pi_{1}(X^{0})\to\pi_{1}(X)$ is surjective with kernel generated by $\delta_{i}$’s, $\exp(2\pi\mathrm{i}\int_{\delta_{i}}\phi_{j})=1$ for all $i$, and $\phi_{j}$ is $d$-closed, the expression $\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})$ is well-defined. Hence, we have (ii) by the representation (1).
(3) By $t_{j}(\gamma x)=\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})t_{j}(x)$ for all $\gamma\in\pi_{1}(X)$, we have $t_{j}$ is a holomorphic nowhere vanishing section of $p^{*}L_{j}$ over $X^{0}$. Let $\{x_{i}\}$ be a coordinate system on an open subset $U$ of $X$ such that $D=(x_{1}x_{2}\cdots x_{c}=0)$. Then
$\phi_{j}|_{U}=\sum_{i=1}^{c}\frac{a_{ij}}{2\pi\mathrm{i}}\frac{dx_{i}}{x_{i}}+\text{some holomorphic function},\text{ and }$
$t_{j}|_{U}=\prod_{i=1}^{c}x_{i}^{a_{ij}}\times\text{some holomorphic function without zeros}.$
Hence $t_{j}$ is a meromorphic, and so rational (by Serre’s GAGA) section of $p^{*}L_{j}$ and $\mathrm{div}(t_{j})=\sum_{i}a_{ij}D_{i}$. ∎
Let $(z_{1},\cdots,z_{d})$ be the inhomogeneous fiber coordinates of the $(\mathbb{P}^{1})^{d}$-bundle $\mathbb{P}_{A}\to A$, the rational map $X\dashrightarrow\mathbb{P}_{A}$ over $A$ is now given by
$x\mapsto(z_{1},\cdots,z_{d})=(t_{1},\cdots,t_{d})\in\mathbb{P}_{A}.$
Let $H:=\mathbb{P}_{A}\setminus G=\sum_{i}((z_{i}=0)+(z_{i}=\infty))$. The pair $(\mathbb{P}_{A},H)$ is log smooth and $K_{\mathbb{P}_{A}}+H\sim 0$.
Example 10.
Let $(X,D_{X})$ be a log smooth pair with $D_{X}$ reduced.
(1) Let $X=\mathbb{P}^{2}$. If $D_{X}=$ 3 lines, then $q(X,D_{X})=3$. If $D_{X}=$ an elliptic curve, then $q(X,D_{X})=0$. If $D_{X}=Q+L$, where $L=(l=0)$ is a line and $Q=(q=0)$ is a conic, then $q(X,D_{X})=1$, $H^{0}(\Omega^{1}_{X}(\log D_{X}))=\mathbb{C}\cdot d\log\frac{l^{2}}{q}$ and the quasi-Albanese morphism is $X\setminus D\to\mathbb{C}^{*},x\mapsto\frac{l^{2}}{q}$.
(2) Let $X=\mathbb{P}^{1}_{z_{1}}\times\mathbb{P}^{1}_{z_{2}}$ and $D_{X}=F_{1}+F_{2}+\Delta$ where $F_{1}=(z_{1}=0)$, $F_{2}=(z_{2}=0)$, $\Delta=(z_{1}=z_{2})$. Then $H^{0}(\Omega_{X}(\log D_{X}))=\mathbb{C}\cdot d\log(1/z_{1}-1/z_{2})$ and the quasi-Albanese morphism is $X\setminus D\to\mathbb{C}^{*},(z_{1},z_{2})\mapsto 1/z_{1}-1/z_{2}$.
(3) Let $A$ be an abelian variety and let $L$ be a line bundle on $A$. Let $X=\mathbb{P}_{A}(O_{A}\oplus L)$, let $z$ be its inhomogeneous fiber coordinate, and let $H_{0}=(z=0)$, $H_{\infty}=(z=\infty)$, $D_{X}=H_{0}+H_{\infty}$. Then $q(X)=\dim A$.
If $L\in\text{Pic}^{0}(A)$, then $q(X,D_{X})=\dim A+1$, and $\mathbb{P}_{A}\setminus H_{1}+H_{2}$ itself is a quasi-abelian variety.
If $L\not\in\text{Pic}^{0}(A)$, then $q(X,D_{X})=\dim A$, and the quasi-Albanese morphism is the natural projection $\mathbb{P}_{A}\setminus H_{1}+H_{2}\to A$.
Remark 11.
Let $X$ be a (possibly) singular variety. Let $\pi:X^{\prime}\to X$ be a resolution and let $f^{\prime}:X^{\prime}\to G$ be the quasi-Albanese morphism of $X^{\prime}$, then the quasi-Albanese map of $X$ is defined as the rational map $f^{\prime}\circ\pi^{-1}$. Assume that $X$ has klt singularities. Then the fibers of $\pi$ are rationally chain connected by Hacon-McKernan [HM07]. Since $G$ contains no rational curves, $f^{\prime}$ factors through $\pi$ by the rigidity Lemma (cf.[Deb01, Lemma 1.15]). The statement in [Deb01, Lemma 1.15] requires $f^{\prime}$ to be proper, but the proof only used that the image of $X^{\prime}\to X\times G$ is proper over $X$, which is guaranteed by [Har77, Chapter II Exercise 4.4].
Proposition 12.
Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $\kappa(K_{X}+D_{X})=0$. Then the quasi-Albanese morphism $X\setminus D_{X}\to G$ is dominant and has irreducible general fibers.
Proof.
Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}:=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Then $q(Y,D_{Y})=q(Y,\pi^{-1}D_{X})=q(X,D_{X})$ and $\kappa(K_{Y}+D_{Y})=0$ by Proposition 5. By the universal property of quasi-Albanese morphisms, we have the following commutative diagram.
[…]
where $g$, $g^{\prime}$ and $f$ are quasi-Albanese morphisms. Since the injective homomorphism $H^{0}(\Omega^{1}_{Y}(\log\pi^{-1}D_{X}))\to H^{0}(\Omega_{Y}^{1}(\log D_{Y}))$ is an isomorphism, we have $H^{1}(Y\setminus D_{Y},\mathbb{C})=H^{1}(O_{Y})\oplus H^{0}(\Omega_{Y}(\log\pi^{-1}D_{X}))\to H^{1}(Y\setminus\pi^{-1}D_{X},\mathbb{C})=H^{1}(O_{Y})\oplus H^{0}(\Omega_{Y}^{1}(\log D_{Y}))$ is an isomorphism by Deligne’s theory of mixed Hodge structures. Hence the surjective morphism $H_{1}(Y\setminus D_{Y},\mathbb{Z})/{\mathrm{torsion}}\to H_{1}(Y\setminus\pi^{-1}D_{X},\mathbb{Z})/{\mathrm{torsion}}$ is also an isomorphism, so $G^{\prime}\to G$ is an isomorphism.
Since $\kappa(K_{Y}+D_{Y})=0$, we have $g^{\prime}$ is dominant and has irreducible general fibers by Theorem 1. Hence, $g$, $f$ are dominant and have irreducible general fibers. ∎
2.3. Subadditivity and positivity
Theorem 13.
Let $f:X\to Y$ be a fibration between smooth projective varieties. Let $D_{X}$ and $D_{Y}$ be reduced snc divisors on $X$ and $Y$ respectively. Denote by $F$ a general fiber of $f$. Assume that $f^{-1}D_{Y}\subset D_{X}$. Then the following holds.
(1)[Mae86, Corollary 2] Let $B$ be a big $\mathbb{Q}$-divisor on $Y$, then
$\kappa(K_{X}+D_{X}-f^{*}(K_{Y}+D_{Y})+f^{*}B)\geq\kappa(K_{F}+{D_{X}}|_{F})+\dim Y.$
(2) If $\kappa(K_{F}+{D_{X}}|_{F})\geq 0$, then $K_{X}+D_{X}-f^{*}(K_{Y}+D_{Y})$ is pseudoeffective.
[…]
Since the $\mathbb{Q}$-divisor $B$ in (1) can be sufficiently small, (2) follows.
Let $X^{0}$ be a smooth quasi-projective variety and let $X^{0}\subset X$ be a compactification such that $D_{X}:=X\setminus X^{0}$ is a simple normal crossing divisor. By the same reasoning as in the definition of $\overline{\kappa}(X^{0})$ (cf. [Fuj24, Lemma 2.3]), we can define the numerical logarithmic Kodaira dimension as $\overline{\nu}(X^{0}):=\nu(K_{X}+D_{X})$. By Fujino’s subadditivity 13(3), we have the following.
Lemma 14.
Let $f:X\to Y$ be a dominant morphism with irreducible general fibers between smooth quasi-projective varieties. Let $F$ be a general fiber of $f$. Then $\overline{\nu}(X)\geq\overline{\nu}(F)+\overline{\kappa}(Y)$.
3. Proof of the main theorems
Proof of Theorem 2.
Let $\alpha:X\setminus D_{X}\to G$ be the quasi-Albanese morphism. Let $G\subset\mathbb{P}_{A}=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1})$ be the compactification described in remark 9, where $L_{i}$’s are elements in $\text{Pic}^{0}(A)$ and let $p:\mathbb{P}_{A}\to A$ be the natural projection. Then $H:=\mathbb{P}_{A}\setminus G$ is a snc divisor and $K_{\mathbb{P}_{A}}+H\sim 0$.
Let $\pi:Z\to X$ be a log resolution of $(X,D_{X})$ resolving the indeterminacies of the rational map $X\dashrightarrow\mathbb{P}_{A}$ and let $D_{Z}:=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. By construction, we have
(3) $f^{-1}H\subset D_{Z}.$
Write $K_{Z}+D_{Z}=\pi^{*}(K_{X}+D_{X})+E_{Z}$. Since $(X,D_{X})$ is lc, $E_{Z}$ is effective and exceptional over $X$. Then by Theorem 15, $(Z,D_{Z})$ has a good minimal model $(W,D_{W})$ over $\mathbb{P}_{A}$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$, the canonical bundle formula for $(W,D_{W})\to\mathbb{P}_{A}$ writes $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H)$, the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, and $g^{-1}H=D_{W}^{v}$.
[…]
Since the birational map $Z\dashrightarrow W$ contracts precisely the divisor $E_{Z}\subset\mathrm{Ex}(\pi)$, we have $\mu:W\dashrightarrow X$ is a birational contraction and $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$. Let $X\xleftarrow{u}V\xrightarrow{v}W$ be a common log resolution of $(X,D_{X})$ and $(W,D_{W})$. Then $u^{*}(K_{X}+D_{X})=v^{*}(K_{W}+D_{W})$ by the Negativity Lemma, i.e., $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ is crepant. By Lemma 6, we have $q(W,D_{W})=q(X,D_{X})$ and $q(W)=q(X)$. By the universal property of quasi-Albanese morphisms and Proposition 12, the composite morphism $W\setminus D_{W}\to W\setminus g^{-1}H\to G$ is the quasi-Albanese morphism of $W\setminus D_{W}$. By the same reasoning, $p\circ g:W\to A$ is the Albanese morphism of $W$. ∎
Theorem 15.
Let $(Z,D_{Z})$ be a projective log smooth pair with $D_{Z}$ reduced and $\kappa(K_{Z}+D_{Z})=0$. Let $A$ be the Albanese variety of $Z$. Let $G$ be a quasi-abelian variety whose abelian part is $A$. Let
$\mathbb{P}_{A}=\mathbb{P}_{A}(O_{A}\oplus L_{1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d})$
be a compactification of $G$ where $L_{i}$’s are elements in $\text{Pic}^{0}(A)$. Let $(z_{1},\cdots,z_{d})$ be the inhomogeneous fiber coordinates of $\mathbb{P}_{A}$. For each $i$, let $p_{i}:\mathbb{P}_{A}\to\mathbb{P}_{A}(O_{A}\oplus L_{i})$ be the projection, and let
$H_{i}=(z_{i}=0)+(z_{i}=\infty)\subset\mathbb{P}_{A}(O_{A}\oplus L_{d})$
be the effective and ample over $A$ divisor. Let $H=\sum_{i}p_{i}^{*}H_{i}$.
Assume that there are
  • a fibration $f:Z\to\mathbb{P}_{A}$ with $f^{-1}H\subset D_{Z}$,
  • a projective $\mathbb{Q}$-factorial dlt pair $(X,D_{X})$ with $K_{X}+D_{X}\sim_{\mathbb{Q}}0$, and
  • a birational morphism $\pi:Z\to X$ such that $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$ and
    $K_{Z}+D_{Z}=\pi^{*}(K_{X}+D_{X})+E_{Z}$
    for some effective and exceptional over $X$ divisor $E_{Z}$.
Then the following holds.
  1. (1)
    Any exceptional over $\mathbb{P}_{A}$ divisor is contained in $E_{Z}+f^{-1}H$.
  2. (2)
    Let $P\subset\mathbb{P}_{A}$ be a prime divisor with $P\not\subset H$, then there is a prime divisor $Q\subset Z$ with $Q\not\subset E_{Z}$ dominating $P$.
  3. (3)
    Assume that for any prime divisor $P\subset H$, there is a prime divisor $Q\not\subset E_{Z}$ dominating $P$, then $(Z,D_{Z})$ has a good minimal model $(W,D_{W})$ over $\mathbb{P}_{A}$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$.
  4. (4)
    Assume that $(Z,D_{Z})$ has a good minimal model $(W,D_{W})$ over $\mathbb{P}_{A}$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$. Let $g:W\to\mathbb{P}_{A}$ be the induced fibration. Then the canonical bundle formula for $(W,D_{W})\to\mathbb{P}_{A}$ has the form
    $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H),$
    the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, and the vertical part $D_{W}^{v}$ of $D_{W}$ coincides with $g^{-1}H$.
  5. (5)
    $(Z,D_{Z})$ has a good minimal model $(W,D_{W})$ over $\mathbb{P}_{A}$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$.
Proof.
Note that by assumption, $G=\mathbb{P}_{A}\setminus H$, $(\mathbb{P}_{A},H)$ is log smooth, and $K_{\mathbb{P}_{A}}+H\sim 0$. Let $F$ be a general fiber of $f$. By Fujino’s subadditivity Theorem 13(3), we have
$0=\nu(K_{Z}+D_{Z})\geq\nu(K_{F}+D_{Z}|_{F})+\kappa(K_{\mathbb{P}_{A}}+H),$
so $\nu(K_{F}+D_{Z}|_{F})=0$. By Gongyo’s abundance theorem [Gon11], $(F,D_{Z}|_{F})$ has a good minimal model and $\kappa(K_{F}+D_{Z}|_{F})=0$.
(1) Let $Q\subset Z$ be an exceptional over $\mathbb{P}_{A}$ prime divisor which is not contained in $f^{-1}H$. Let $q:\mathbb{P}^{\prime}_{A}\to\mathbb{P}_{A}$ be a composition of the blowup along $f(Q)$ with a log resolution. We may assume that $q$ is an isomorphism over $\mathbb{P}_{A}\setminus f(Q)$. Let $p:Z^{\prime}\to Z$ be a log resolution of $(Z,D_{Z})$ resolving the indeterminacies of $Z\dashrightarrow\mathbb{P}_{A}^{\prime}$. We have the following commutative diagram.
[…]
Write $K_{Z^{\prime}}+D_{Z^{\prime}}=p^{*}(K_{Z}+D_{Z})+E_{1}$, $K_{\mathbb{P}^{\prime}_{A}}+H^{\prime}=q^{*}(K_{\mathbb{P}_{A}}+H)+E_{2}$, where $H^{\prime}=q^{-1}H$, $D_{Z^{\prime}}=p^{-1}_{*}D_{Z}+\operatorname{Ex}(p)$, and $E_{1}$, $E_{2}$ are effective and exceptional divisors. By $f^{\prime-1}H^{\prime}\subset D_{Z^{\prime}}$,
$0\sim_{\mathbb{Q}}p^{*}(K_{Z}+D_{Z}-f^{*}(K_{\mathbb{P}_{A}}+H)-E_{Z})=K_{Z^{\prime}}+D_{Z^{\prime}}-f^{\prime*}(K_{\mathbb{P}^{\prime}_{A}}+H^{\prime})-E_{1}+f^{\prime*}E_{2}-p^{*}E_{Z},$
and Theorem 13(1), we have $E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z}$ is pseudoeffective. Since $(\pi\circ p)_{*}(E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z})=-(\pi\circ p)_{*}f^{\prime*}E_{2}$ is pseudoeffective, we have $f^{\prime*}E_{2}$ is exceptional over $X$. Hence $E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z}$ is effective by a Lemma due to Lazarsfeld (cf. [KL09, Corollary 13]).
Let $Q^{\prime}$ be the strict transform of $Q$ on $Z^{\prime}$ and let $P^{\prime}=f^{\prime}(Q^{\prime})$. Since $q(P^{\prime})=f(Q)\not\subset H$, we have $\operatorname{mult}_{P^{\prime}}E_{2}>0$, so $\operatorname{mult}_{Q^{\prime}}f^{\prime*}E_{2}>0$. By
$0\leq\operatorname{mult}_{Q^{\prime}}(E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z})=\operatorname{mult}_{Q^{\prime}}(-f^{\prime*}E_{2}+p^{*}E_{Z}),$
we have $Q^{\prime}\subset p^{*}E_{Z}$. Hence $Q\subset E_{Z}$.
(2) Assume the contrary, i.e., assume that
$f^{-1}P\subset E_{Z}+E$
for some effective exceptional over $\mathbb{P}_{A}$ divisor $E$. Then by (1), we have
$f^{-1}(P+H)\subset E_{Z}+f^{-1}H\subset D_{Z}.$
Then by Lemma 14 and Lemma 16, we have
$0=\nu(K_{Z}+D_{Z})=\overline{\nu}(Z\setminus D_{Z})\geq\overline{\nu}(F\setminus D_{Z})+\overline{\kappa}({\mathbb{P}_{A}}\setminus P+H)=\overline{\kappa}(G\setminus P)>0,$
which is a contradiction.
(3) By (2) and the assumption in (3), we known that the vertical part $E_{Z}^{v}$ of $E_{Z}$ is very exceptional over $\mathbb{P}_{A}$. Since $\kappa(E_{Z}|_{F})=\kappa(K_{F}+D_{Z}|_{F})=0$, we have
(4) $(f_{*}O_{Z}(mE_{Z}))^{\wedge}=O_{\mathbb{P}_{A}}$
for all sufficiently divisible $m>0$, where $\wedge$ means taking reflexive hull.
By Gongyo’s abundance theorem and Hacon-Xu [HX13, Theorem 2.12], $(Z,D_{Z})$ has a good minimal model over a nonempty Zariski open subset $U$ of $\mathbb{P}_{A}$. We run a $(Z,D_{Z})$-MMP over $\mathbb{P}_{A}$ with scaling of an ample divisor. Then this MMP will terminate over $U$ by [HX13, Corollary 2.9]. Hence we have a partial MMP $(Z,D_{Z})\dashrightarrow(Z^{\prime},D_{Z^{\prime}})$ with $K_{Z^{\prime}}+D_{Z^{\prime}}$ semiample over $U$, where $D_{Z^{\prime}}$ is the pushforward of $D_{Z}$. Let $F^{\prime}$ be a general fiber of $f^{\prime}:Z^{\prime}\to\mathbb{P}_{A}$ and let $E_{Z^{\prime}}$ be the pushforward of $E_{Z}$. Since $\kappa(K_{F^{\prime}}+D_{Z^{\prime}}|_{F^{\prime}})=0$, we have $E_{Z^{\prime}}|_{F^{\prime}}\sim_{\mathbb{Q}}K_{Z^{\prime}}+D_{Z^{\prime}}|_{F^{\prime}}\sim_{\mathbb{Q}}0$, so $E_{Z^{\prime}}$ is vertical over $\mathbb{P}_{A}$. By formula (4), we have $(f^{\prime}_{*}O_{Z^{\prime}}(mE_{Z^{\prime}}))^{\wedge}=(f_{*}O_{Z}(mE_{Z}))^{\wedge}=O_{\mathbb{P}_{A}}$ for all sufficiently divisible $m>0$. Hence $E_{Z^{\prime}}$ is very exceptional very $\mathbb{P}_{A}$ by [GL13, Lemma 2.15]. Then $(Z^{\prime},D_{Z^{\prime}})$, and so $(Z,D_{Z})$, has a good minimal model over $\mathbb{P}_{A}$ by Birkar [Bir12, Theorem 1.8].
(4) Write the canonical bundle formula for $g:(W,D_{W})\to\mathbb{P}_{A}$ as $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+B_{\mathbb{P}_{A}}+M_{\mathbb{P}_{A}})$. Since $f^{-1}H\subset D_{Z}$, we have $g^{-1}H\subset D_{W}$, so $B_{\mathbb{P}_{A}}\geq H$. Since $K_{W}+D_{W}\sim_{\mathbb{Q}}0$ and $K_{\mathbb{P}_{A}}+H\sim 0$, we have $B_{\mathbb{P}_{A}}=H$ and $M_{\mathbb{P}_{A}}\sim_{\mathbb{Q}}0$, so the canonical bundle formula writes
(5) $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H).$
Let $\Sigma$ be a reduced snc divisor on $G$ and let $\overline{\Sigma}$ be its closure in $\mathbb{P}_{A}$. Then the canonical bundle formula for $(W,D_{W}+g^{*}\overline{\Sigma})\to\mathbb{P}_{A}$ writes $K_{W}+D_{W}+g^{*}\overline{\Sigma}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H+\overline{\Sigma})$. By inversion of adjunction (cf. [Fil20, Proposition 4.16]), $(W,D_{W}+g^{*}\overline{\Sigma})|_{g^{-1}G}$ is lc. Hence the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable by [Kol23, Corollary 4.55]. In particular, the restriction $g|_{g^{-1}G}$ is equidimensional, and so $W$ has no exceptional divisors over $G$.
Let $Q\subset D_{W}^{v}$ be a prime divisor. If $g(Q)\not\subset H$, then $g(Q)$ is a divisor on $\mathbb{P}_{A}$. But by the canonical bundle formula (5), we must have $H\geq g(Q)$, which is a contradiction. Hence $D_{W}^{v}=g^{-1}H$.
(5) By (3), it suffices to show that for any prime divisor $P\subset H$, there is a prime divisor $Q\not\subset E_{Z}$ dominating $P$. Assume the contrary, i.e., assume that there is a prime divisor $P\subset H$, and an effective exceptional over $\mathbb{P}_{A}$ divisor $E$ with $f(E)\subset P$ such that
(6) $f^{-1}P\subset E_{Z}+E.$
Case 1. $\dim\mathbb{P}_{A}-\dim A=1$. By [BFPT24, Theorem 1.5], the Albanese morphism $p_{X}:X\to A^{\prime}$ of $X$ is a fibration. Since $X$ has rational singularities, we have $\dim A^{\prime}=q(X)=q(Z)=\dim A$. Hence, the induced morphism $A\to A^{\prime}$ is an isomorphism. We identify $A^{\prime}$ with $A$. We have the following commutative diagram.
[…]
where $p$ is the natural projection.
We claim that
(7) $f^{-1}P\subset E_{Z}.$
Let $Q\subset E$ be a prime divisor. Then $Q\subset f^{-1}H\subset D_{Z}$. Since $\dim\mathbb{P}_{A}-\dim A=1$, $f(Q)$ is vertical over $A$, so $f(Q)$ is not an lc center of $(X,D_{X})$ by [BFPT24, Theorem 1.5]. Hence
$1<a(Q,X,D_{X})=-\operatorname{mult}_{Q}(D_{Z}-E_{Z})=-1+\operatorname{mult}_{Q}E_{Z}.$
Hence $Q\subset E_{Z}$ and the claim follows.
Let $L_{A}$ be an effective and sufficiently ample divisor on $A$. Since $P$ is ample over $A$, we have
$\dim\mathbb{P}_{A}$ $=\kappa(P+p^{*}L_{A})$
$=\kappa(f^{*}P+(p\circ f)^{*}L_{A})$
[…]
$=\kappa(E_{Z}+(p_{X}\circ\pi)^{*}L_{A})$
$=\kappa(p_{X}^{*}L_{A})\text{ ($E_{Z}$ is exceptional over $X$)}$
$=\dim A,$
which is a contradiction, so the inclusion (6) can not be true.
Case 2. $\dim\mathbb{P}_{A}-\dim A>1$. Since $P\subset H=\sum_{i}p_{i}^{*}H_{i}$, we have $P\subset p_{i}^{*}H_{i}$ for some $i$. By re-indexing $i$’s, we may assume that $P\subset p_{1}^{*}H_{1}$. Let
$\mathbb{P}^{\prime}_{A}:=\mathbb{P}_{A}(O_{A}\oplus L_{2})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}),$
let $p_{i}^{\prime}:\mathbb{P}_{A}^{\prime}\to\mathbb{P}_{A}(O_{A}\oplus L_{i})$, $2\leq i\leq d$, $\alpha:\mathbb{P}_{A}\to\mathbb{P}_{A}^{\prime}$ be the natural projections and let $H^{\prime}=\sum_{i=2}^{d}{p^{\prime}}_{i}^{*}H_{i}$. Then
$(\alpha\circ f)^{-1}H^{\prime}\subset f^{-1}H\subset D_{Z}.$
By induction on $(\dim\mathbb{P}_{A}-\dim A)$, $(Z,D_{Z})$ has a good minimal model $(V,D_{V})$ over $\mathbb{P}_{A}^{\prime}$ with $K_{V}+D_{V}\sim_{\mathbb{Q}}0$. We have the following commutative diagram.
[…]
where $p^{\prime}$ is the natural projection and $p_{V}$ is the Albanese morphism of $V$ (by the same reasoning as in Case 1, we may identify the Albanese varieties of $Z$ and $V$).
We claim that
(8) $f^{-1}P\subset E_{Z}+(\alpha\circ f)^{-1}H^{\prime}.$
Let $Q\subset E$ be a prime divisor which is not contained in $E_{Z}$. Then $Q\subset f^{-1}H\subset D_{Z}$. Since $Z\dashrightarrow V$ contracts precisely the divisor $E_{Z}$, the strict transform $Q_{V}$ of $Q$ on $V$ is a divisor contained in $D_{V}$. Since $\dim\mathbb{P}_{A}-\dim\mathbb{P}_{A}^{\prime}=1$, and $Q$ is exceptional over $\mathbb{P}_{A}$, we have $Q$ is vertical over $\mathbb{P}_{A}^{\prime}$. Hence $Q_{V}$ is vertical over $\mathbb{P}_{A}^{\prime}$. By (4), we have $g(Q_{V})\subset H^{\prime}$. Hence $Q\subset(\alpha\circ f)^{-1}H^{\prime}$ and the claim follows.
Let $L_{A}$ be an effective and sufficiently ample divisor on $A$. Since $H^{\prime}$ and $\alpha^{*}H^{\prime}+P$ are ample over $A$, we have
$\dim\mathbb{P}_{A}$ $=\kappa(P+\alpha^{*}H^{\prime}+(p^{\prime}\circ\alpha)^{*}L_{A})$
$=\kappa(f^{*}P+(\alpha\circ f)^{*}H^{\prime}+(p^{\prime}\circ\alpha\circ f)^{*}L_{A})$
[…]
$=\dim\mathbb{P}_{A}^{\prime},$
which is a contradiction, so the inclusion (6) can not be true. ∎
Lemma 16.
Let $G$ be a quasi-abelian variety and let $P\subset G$ be a prime divisor, then $\overline{\kappa}(G\setminus P)>0$.
Proof.
The idea is from [Fuj24, Proof of Theorem 1.3, Step 2]. Let $A$ be the abelian part of $G$ and let $p:G\to A$ be the natural projection.
Case 1. $P$ is vertical over $A$. Then $P=p^{*}Q$ for some prime divisor $Q$ on $A$. Let $\mathbb{P}_{A}\to A$ be a $(\mathbb{P}_{1})^{d}$-bundle over $A$ such that $\mathbb{P}_{A}$ is a compactification of $G$, and $H=\mathbb{P}_{A}\setminus G$ is a snc divisor with $K_{\mathbb{P}_{A}}+H\sim 0$. Let $\overline{P}$ be the closure of $P$ in $\mathbb{P}_{A}$. Since any lc center of $(\mathbb{P}_{A},H)$ is horizontal over $A$, $\overline{P}$ contains no lc centers of $(\mathbb{P}_{A},H)$. Then $\overline{\kappa}(G\setminus P)=\kappa(K_{\mathbb{P}_{A}}+H+\overline{P})=\kappa(\overline{P})=\kappa(Q)>0$ by Lemma 17.
Case 2. $P$ is horizontal over $A$. Then there is a subgroup $\mathbb{C}^{*}\subset G$ such that $P$ dominates the quasi-abelian variety $G^{\prime}:=G/\mathbb{C}^{*}$. Let $F$ be a general fiber of $G\setminus P\to G^{\prime}$. Then $F$ is a curve with $\overline{\kappa}(F)>0$. Hence $\overline{\kappa}(G\setminus P)\geq\overline{\kappa}(F)+\overline{\kappa}(G^{\prime})=\overline{\kappa}(F)>0$ by Kawamata [Kaw78, Theorem 1] (see also Fujino [Fuj20, Chapter 5]).∎
Lemma 17.
Let $(Y,D_{Y})$ be a projective log smooth pair with $D_{Y}$ reduced. Assume that $\kappa(K_{Y}+D_{Y})\geq 0$. Let $\Sigma\subset Y$ be a reduced divisor which contains no lc centers of $(Y,D_{Y})$, then $\overline{\kappa}(Y\setminus D_{Y}+\Sigma)=\kappa(K_{Y}+D_{Y}+\Sigma)$.
Proof.
Let $\mu:Z\to Y$ be a log resolution of $(Y,D_{Y}+\Sigma)$. We may assume that $\mu$ is an isomorphism over $Y\setminus D_{Y}+\Sigma$. Let $\epsilon>0$ be a sufficiently small rational number. Then
$\overline{\kappa}(Y\setminus D_{Y}+\Sigma)$ $=\overline{\kappa}(Z\setminus\mu^{-1}(D_{Y}+\Sigma))$
$=\kappa(K_{Z}+\mu^{-1}(D_{Y}+\Sigma))$
$=\kappa(K_{Z}+\mu^{-1}_{*}D_{Y}+\mu^{-1}_{*}\Sigma+\operatorname{Ex}(\mu))$
$=\kappa(K_{Z}+\mu^{-1}_{*}D_{Y}+\epsilon\mu^{-1}_{*}\Sigma+\operatorname{Ex}(\mu))\text{ (by $\kappa(K_{Z}+\mu^{-1}_{*}D_{Y}+\operatorname{Ex}(\mu))\geq 0$)}$
$=\kappa(K_{Y}+D_{Y}+\epsilon\Sigma)\text{ (by $(Y,D_{Y}+\epsilon\Sigma)$ is lc)}.$
$=\kappa(K_{Y}+D_{Y}+\Sigma)\text{ (by $\kappa(K_{Y}+D_{Y})\geq 0$)}.$
Lemma 18.
Let $\alpha:Z\to V$ and $\beta:W\to V$ be fibrations between projective normal varieties, and let $Z\dashrightarrow W$ be a birational contraction over $V$. Let $E$ be an effective $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor on $Z$ which is exceptional over $W$, and let $L$ be a $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor on $V$. Then $\kappa(\alpha^{*}L+E)=\kappa(L)$.
Proof.
Let $u:X\to Z$ be a birational morphism resolving the indeterminacies of $Z\dashrightarrow W$ and let $v:X\to W$ be the induced morphism. Since $Z\dashrightarrow W$ is a birational contraction, $u^{*}E$ is exceptional over $W$. Hence
$\kappa(\alpha^{*}L+E)=\kappa(u^{*}\alpha^{*}L+u^{*}E)=\kappa(v^{*}\beta^{*}L+u^{*}E)=\kappa(\beta^{*}L)=\kappa(L).$
Theorem 2 is a special case of the following theorem.
Theorem 19.
Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$. Let $\alpha:X\setminus D\to G$ be the quasi-Albanese morphism. Then the conclusion in Theorem 2 holds.
Proof.
Keep the notations introduced in the proof of Theorem 3.
(1) Let $P\subset G$ be a prime divisor and let $\overline{P}$ be its closure in $\mathbb{P}_{A}$. Then $\overline{P}\not\subset H$. Since $D_{W}^{v}=g^{-1}H$, there is a prime divisor $\overline{Q}_{W}\subset W$ such that $\overline{Q}_{W}\not\subset D_{W}$ and $\overline{Q}_{W}$ dominates $\overline{P}$. Let $\overline{Q}$ be the strict transform of $\overline{Q}_{W}$ on $X$. Then $\overline{Q}\not\subset D_{X}$ and $\overline{Q}$ dominates $P$.
(2) Let $Q\subset X\setminus D_{X}$ be a prime divisor and let $\overline{Q}$ be its closure in $X$. Then $\overline{Q}\not\subset D_{X}$ and its strict transform $\overline{Q}_{W}$ on $W$ has $\overline{Q}_{W}\not\subset D_{W}$, so $\overline{Q}_{W}\not\subset g^{-1}H$. Since $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable and in particular equidimensional, we have $\operatorname{codim}\alpha(Q)=\operatorname{codim}g(\overline{Q}_{W})=1$.
(3) Let $P\subset G$ be a prime divisor and let $\overline{P}$ be its closure in $\mathbb{P}_{A}$. Since $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, we have $(W,D_{W}+g^{*}\overline{P})$ is lc near the generic point of $\overline{P}$, so $(Z,D_{Z}-E_{Z}+f^{*}\overline{P})|_{f^{-1}G}$ is sub-lc near the generic point of $P$. By
$(K_{Z}+D_{Z}-E_{Z}+f^{*}\overline{P})|_{Z\setminus\pi^{-1}D_{X}}=(\pi|_{Z\setminus\pi^{-1}D_{X}})^{*}(K_{X\setminus D_{X}}+\alpha^{*}P)$
and $Z\setminus\pi^{-1}D_{X}\subset Z\setminus f^{-1}H=f^{-1}{G}$, we have $(X\setminus D_{X},\alpha^{*}P)$ is lc near the generic point of $P$. ∎
4. A remark on orbifold fundamental groups
This section is inspired by the work [GLM23]. Let $(X,D_{X}=\sum_{i}(1-\frac{1}{m_{i}})D_{i})$ be an lc pair with standard coefficients, where each $D_{i}$ is a prime divisor and $m_{i}\in\mathbb{Z}_{>0}\cup\infty$. For each $i$, let $\gamma_{i}$ be a small loop around $D_{i}$. The orbifold fundamental group $\pi_{1}^{\operatorname{orb}}(X,D_{X})$ is defined as the quotient of $\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})$ by the normal subgroup generated by $\gamma_{i}^{m_{i}}$’s. Note that if $D_{X}$ is reduced, then $\pi_{1}^{\operatorname{orb}}(X,D_{X})=\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})$.
Conjecture 20.
Let $(X,D_{X})$ be an lc pair with standard coefficients. If $K_{X}+D_{X}\sim_{\mathbb{Q}}0$, then $\pi_{1}^{\operatorname{orb}}(X,D)$ is virtually nilpotent.
Using the index-one cover and the Galois correspondence for orbifold fundamental groups (cf. [GLM23, Proposition A.11]), Conjecture 20 is reduced to the case $D_{X}$ is reduced. By the following Lemma, Conjecture 20 is further reduced to the $\mathbb{Q}$-factorial dlt case.
Lemma 21.
Let $(X,D_{X})$ and $(Y,D_{Y})$ be two projective lc pairs with $D_{X}$ and $D_{Y}$ reduced. If there is a birational contraction $p:X\dashrightarrow Y$ such that $D_{X}=p^{-1}_{*}D_{Y}+\operatorname{Ex}(p)$, then there is an isomorphism $\pi_{1}^{\operatorname{orb}}(X,D_{X})\cong\pi_{1}^{\operatorname{orb}}(Y,D_{Y})$.
Proof.
Let $Z\subset X$ and $Z^{\prime}\subset Y$ be closed subsets such that $X\setminus Z\cong Y\setminus Z^{\prime}$. Hence $X^{\operatorname{reg}}\setminus Z\cong Y^{\operatorname{reg}}\setminus Z^{\prime}$ and $X^{\operatorname{reg}}\setminus D_{X}\cup Z\cong Y^{\operatorname{reg}}\setminus D_{Y}\cup Z^{\prime}$. Since $p$ is a birational contraction, we have $\operatorname{codim}Z^{\prime}\geq 2$ and $D_{X}\cup Z=D_{X}\cup Z^{\prime\prime}$ for some $\operatorname{codim}\geq 2$ closed subset $Z^{\prime\prime}$. Hence $\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})\cong\pi_{1}(X^{\operatorname{reg}}\setminus D_{X}\cup Z^{\prime\prime})\cong\pi_{1}(X^{\operatorname{reg}}\setminus D_{X}\cup Z)\cong\pi_{1}(Y^{\operatorname{reg}}\setminus D_{Y}\cup Z^{\prime})\cong\pi_{1}(Y^{\operatorname{reg}}\setminus D_{Y})$. ∎
Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$.
By a theorem due to Arapura and Nori [AN99, Corollary 4.10], to show that $\pi_{1}^{\text{orb}}(X,D_{X})=\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})$ is virtually nilpotent, it suffices to show that it is polycyclic. Recall that a group $G$ is called polycyclic if there is a series of subgroups $1=G_{n}\triangleleft G_{n-1}\triangleleft\cdots\triangleleft G_{0}=G$ such that each $G_{i}/G_{i+1}$ is cyclic. The advantage of the class of polycyclic groups is that they are closed under passing to subgroups, quotients, and extensions (cf. [KM79, 4.4.3. Exercise]), which is crucial to inductive arguments.
Assume that $q(X,D_{X})>0$, and let $g:(W,D_{W})\to\mathbb{P}_{A}$ be the fibration given by Theorem 3. By Lemma 21, $(X,D_{X})$ and $(W,D_{W})$ have the same orbifold fundamental group. Let $W^{0}=g^{-1}G$ and let $F$ be a general fiber of $g$. Then $(F,{D_{W}}|_{F})$ is lc and $K_{F}+D_{W}|_{F}\sim_{\mathbb{Q}}0$. By the following exact sequence (cf. [GLM23, Lemma 2.18])
$\pi_{1}^{\operatorname{orb}}(F,D_{W}|_{F})\to\pi_{1}^{\operatorname{orb}}(W^{0},{D_{W}}|_{W^{0}})\to\pi_{1}(G)\to 1,$
we have $\pi_{1}^{\operatorname{orb}}(X,D_{X})\cong\pi_{1}^{\operatorname{orb}}(W,D_{W})\cong\pi_{1}^{\operatorname{orb}}(W^{0},{D_{W}}|_{W^{0}})$ is polycyclic if $\pi_{1}^{\operatorname{orb}}(F,D_{W}|_{F})$ is polycyclic. Hence, we have shown the following by induction on dimension.
Proposition 22.
To prove Conjecture 20, it suffices to prove $\mathbb{Q}$-factorial dlt pairs $(X,D_{X})$ with $D_{X}$ reduced and $q(X,D_{X})=0$ have polycyclic orbifold fundamental groups.

### v2

@@
 Abstract.
-We study the quasi-Albanese morphisms for log canonical Calabi-Yau pairs.
+In this note, we study quasi-Albanese morphisms for log canonical Calabi-Yau pairs and obtain several structural results. As an application, we prove a characterization of toric pairs.
 2020 Mathematics Subject Classification:
@@
 1. Introduction
-Let $X$ be a complex projective variety with canonical singularities and $K_{X}$ torsion.
-Kawamata proved that its Albanese morphism $\alpha:X\to A$ is an étale fiber bundle [Kaw85, Theorem 8.3].
-Inspired by the works [Iit76, Kaw81, Zha05, Fuj24, BFPT24], we study quasi-Albanese morphisms for (the open part of) log canonical Calabi-Yau pairs.
-Before stating the main theorem, we recall the definition of logarithmic Kodaira dimension and a theorem by Kawamata and Fujino.
-Let $X^{0}$ be a smooth quasi-projective variety.
-Let $X$ be a compactification of $X^{0}$ such that $D_{X}:=X\setminus X^{0}$ is a simple normal crossing divisor.
-The logarithmic Kodaira dimension $\overline{\kappa}(X^{0})$ of $X^{0}$ is defined as $\kappa(K_{X}+D_{X})$.
-By Iitaka [Iit77], $\overline{\kappa}(X^{0})$ is independent of the choice of such $X$.
-Theorem 1.
-[Kaw81, Thoerem 28], see also [Fuj24, Theorem 1.2] If $\overline{\kappa}(X^{0})=0$, then the quasi-Albanese morphism of $X^{0}$ is dominant and has irreducible general fibers.
-The following is our main theorem.
-Theorem 2.
-Let $(X,D_{X})$ be a log smooth pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$. Let $X^{0}=X\setminus D_{X}$ and let $\alpha:X^{0}\to G$ be its quasi-Albanese morphism. Then
+Let $X$ be a smooth complex projective Calabi-Yau variety.
+By the Bogomolov-Beauville-Yau decomposition, the Albanese morphism of $X$ is isotrivial.
+Kawamata [kawamata1985minimal, Theorem 8.3] extended this isotriviality to varieties with canonical singularities, and Ambro [Ambro2005moduli, Theorem 4.8] to klt pairs.
+For log canonical Calabi-Yau pairs, Bernasconi, Filipazzi, Patakfalvi and Tsakanikas [bernasconi2024strong, Theorem 1.5] proved that the Albanese morphism is locally stable in the sense of Kollár [Kollar-Families, Definition-Theorem 4.7], but may not be isotrivial.
+In the setting of log pairs, it is more natural to consider logarithmic differential forms and quasi-Albanese morphisms by Iitaka [iitaka1976logarithmic], which capture finer structural information than the Albanese morphisms.
+In this note, we study quasi-Albanese morphisms for log canonical Calabi-Yau pairs. Our results are intended as a step toward a log canonical Bogomolov–Beauville–Yau decomposition.
+We briefly recall the relevant definitions. A quasi-abelian variety $G$ is a quasi-projective commutative algebraic group fitting into an exact sequence
+$0\to(\mathbb{C}^{*})^{d}\to G\to A\to 0,$
+where $A$ is an abelian variety and $(\mathbb{C}^{*})^{d}$ is an algebraic torus; we refer to $A$ as the compact part of $G$.
+Let $X$ be a quasi-projective variety.
+The quasi-Albanese morphism $\alpha:X\to G$ is a morphism to a quasi-abelian variety characterized by the following universal property: any morphism from $X$ to a quasi-abelian variety factors uniquely through $\alpha$.
+The variety $G$ is called the quasi-Albanese variety of $X$.
+We recall the following classical result by Kawamata.
+Theorem 1.1 ([kawamata1981characterization, Theorem 28], see also [fujino2024quasi, Theorem 1.2]).
+Let $(X,D_{X})$ be a projective log smooth pair with $D_{X}$ reduced and $\kappa(K_{X}+D_{X})=0$.
+Then the quasi-Albanese morphism $\alpha:X\setminus D_{X}\to G$ is dominant and has irreducible general fibers.
+We now turn to the singular setting. For a log canonical pair $(X,D_{X})$, we denote by $\operatorname{nklt}(X,D_{X})$ its non-klt locus. We strengthen Kawamata’ Theorem as follows.
+Theorem 1.2.
+Let $(X,D_{X})$ be a projective log canonical pair with $\kappa(K_{X}+D_{X})=0$.
+Then the quasi-Albanese morphism $\alpha:X\setminus\operatorname{nklt}(X,D_{X})\to G$ is dominant and has irreducible general fibers.
+The next theorem further analyzes the quasi-Albanese morphism $\alpha$ in the Calabi-Yau case.
+Theorem 1.3.
+Let $(X,D_{X})$ be a projective log canonical pair with $K_{X}+D_{X}\sim_{\mathbb{Q}}0$.
+Let $\alpha:X\setminus\operatorname{nklt}(X,D_{X})\to G$ be the quasi-Albanese morphism.
+Then the following holds.
   1. (1)
-    $\alpha$ is surjective in codimension one: for any prime divisor $P\subset G$, there is a prime divisor $Q\subset X^{0}$ dominating $P$,
+    $\alpha$ is surjective in codimension one: for any prime divisor $P\subset G$, there is a prime divisor $Q\subset X^{0}$ dominating $P$.
   2. (2)
-    $\alpha$ is flat in codimension one: for any prime divisor $Q\subset X^{0}$, if $Q$ do not dominate $G$, then $\alpha(Q)$ is of codimension one,
+    $\alpha:(X\setminus\operatorname{nklt}(X,D_{X}),D_{X}|_{X\setminus\operatorname{nklt}(X,D_{X})})\to G$ is locally stable.
+In particular, $\alpha$ is flat and every fiber is reduced with semi-log canonical singularities.
+Let $G$ be a quasi-abelian variety and let $A$ be its compact part.
+Then there are $d:=\dim G-\dim A$ line bundles $L_{i}\in\text{Pic}^{0}(A)$ such that $G$ is the fiber product over $A$ of the total spaces of the $L_{i}$ with their zero sections removed.
+Consequently, $G$ admits the following natural compactification:
+$G\subset\mathbb{P}_{A}:=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1}).$
+Note that $H:=\mathbb{P}_{A}\setminus G$ is a simple normal crossing divisor, and the log smooth pair $(\mathbb{P}_{A},H)$ is Calabi-Yau: $K_{\mathbb{P}_{A}}+H\sim 0$.
+We refer to $(\mathbb{P}_{A},H)$ as the canonical compactification of $G$.
+Let $(X,D_{X})$ be a log canonical pair.
+Via the quasi-Albanese morphism $\alpha:X\setminus\operatorname{nklt}(X,D_{X})\to G$, we aim to extract structural information about the pair $(X,D_{X})$ itself, rather than merely about the open subset $X\setminus\operatorname{nklt}(X,D_{X})$.
+To this end, we consider the canonical compactification $(\mathbb{P}_{A},H)$ of $G$ and the induced rational map $X\dashrightarrow\mathbb{P}_{A}$.
+In general, however, this rational map need not be a morphism (see Example 2.5).
+The following theorem asserts that for a Calabi-Yau pair $(X,D_{X})$, one can always find another Calabi-Yau pair $(W,D_{W})$, crepant birational to $(X,D_{X})$, such that the induced map $W\dashrightarrow\mathbb{P}_{A}$ is a morphism.
+Theorem 1.4.
+Let $(X,D_{X})$ be a projective log canonical pair with $K_{X}+D_{X}\sim_{\mathbb{Q}}0$.
+Let $\alpha:X\setminus\operatorname{nklt}(X,D_{X})\to G$ be the quasi-Albanese morphism.
+Let $(\mathbb{P}_{A},H)$ be the canonical compactification of $G$.
+Then there are
+  • a projective $\mathbb{Q}$-factorial dlt pair $(W,D_{W})$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$, and
+  • a crepant birational contraction $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ such that $D_{W}=\mu^{-1}_{*}D_{X}+\sum P$, where the sum runs over all prime divisors on $W$ that are exceptional over $X$,
+satisfying the following properties:
   1. (1)
-    $\alpha$ is semistable in codimension one: for any prime divisor $P\subset G$, the pair $(X,\alpha^{*}P)$ is log canonical near the generic point of $P$.
-Theorem 2 is a consequence of the following theorem, whose proof depends on Theorem 1 and the techniques from the minimal model program.
-Theorem 3.
-Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$.
-Set $q(X):=\dim H^{1}(O_{X})$ and $q(X,D_{X}):=\dim H^{0}(\Omega_{X^{0}}^{1}(\log D_{X}|_{X^{0}}))$, where $X^{0}$ is the log smooth locus of $(X,D_{X})$.
-Let $A$ be the Albanese variety of $X$ and let $G$ be the quasi-Albanese variety of $X\setminus D_{X}$.
-Then there are
-  • a $\mathbb{Q}$-factorial dlt pair $(W,D_{W})$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$,
-  • a crepant birational contraction $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ with $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$,
-  • $(q(X,D_{X})-q(X))$’s line bundles $L_{i}\in\mathrm{Pic}^{0}(A)$, and
-  • a commutative diagram,
-[…]
-    where $\mathbb{P}_{A}:=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1})$ is a compactification of $G$, $H=\mathbb{P}_{A}\setminus G$ is a simple normal crossing divisor, $g$ is a fibration, and $p$ is the natural projection,
-such that the following holds.
+    The induced rational map $g:W\dashrightarrow\mathbb{P}_{A}$ is a morphism, which is surjective and has connected fibers.
   2. (2)
-    $p\circ g$ is the Albanese morphism of $W$, $h$ is the quasi-Albanese morphism of $W\setminus D_{W}$.
+    The canonical bundle formula for $g:(W,D_{W})\to\mathbb{P}_{A}$ takes the form
+    $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H),$
+    and the restriction $(W,D_{W})|_{g^{-1}(G)}\to G$ is locally stable.
   3. (3)
-    The $g$-vertical part of $D_{W}$ coincides with $g^{-1}H$.
+    The support of $g^{*}H$ is contained in the non-klt locus $\operatorname{nklt}(W,D_{W})$ and the induced morphism
+    $\beta:W\setminus\operatorname{nklt}(W,D_{W})\to G$
+    is the quasi-Albanese morphism of $W\setminus\operatorname{nklt}(W,D_{W})$.
+As a byproduct of our study, we find a connection between the Shokurov complexity of log canonical pairs and their quasi-Albanese varieties.
+For a log canonical pair $(X,D_{X})$, its complexity $c(X,D_{X})$ is defined as
+$c(X,D_{X}):=\dim X+\operatorname{rank}\langle\lfloor D_{X}\rfloor\rangle-|\lfloor D_{X}\rfloor|,$
+where
+  • $\langle\lfloor D_{X}\rfloor\rangle$ is the subgroup generated by the components of $\lfloor D_{X}\rfloor$ in the group of Weil divisors modulo algebraic equivalence, and
+  • $|\lfloor D_{X}\rfloor|$ is the sum of the coefficients of $\lfloor D_{X}\rfloor$.
+For simplicity, we adopt a definition slightly different from that in [Brown2018GeomToric].
+In that work, Brown, McKernan, Svaldi, and Zong proved a conjecture of Shokurov [Shokurov2000CompleSurf] characterizing toric varieties via complexities.
+See also [Prokhorov2001Toric, yao2013thesis, Nakayama2017Shokurov, enwright2024log] for related results.
+In this note, we show the following.
+Lemma 1.5.
+Let $(X,D_{X})$ be a projective log canonical pair. Let $G$ be the quasi-Albanese variety of $X\setminus\operatorname{nklt}(X,D_{X})$, and let $A$ be its compact part. Then
+$c(X,D_{X})=\dim X-(\dim G-\dim A).$
+Theorem 1.6.
+Let $(X,D_{X})$ be a projective log canonical pair with $\kappa(K_{X}+D_{X})=0$. Then, for any reduced divisor $0\leq\Delta_{X}\leq\lfloor D_{X}\rfloor$, we have
+$c(X,\Delta_{X})\geq 0.$
+Assume that the equality holds for some $\Delta_{X}$. Then $X$ is rational. Assume further that $K_{X}+D_{X}\sim_{\mathbb{Q}}0$, then the following holds.
   1. (1)
-    The canonical bundle formula for $g:(W,D_{W})\to\mathbb{P}_{A}$ has the form
-    $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H),$
-    and the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable.
-As a corollary, if $q(X)=0$ and $q(X,D_{X})=\dim X$, then $(X,D_{X})$ is crepant birational to $(\mathbb{P}^{n},(\prod_{i}z_{i}=0))$.
-Another corollary is that if $q(X,D_{X})-q(X)\geq 2$, then $D_{X}$ is connected.
-In the previous Theorem, the pair $(W,D_{W})$ is closely related to $(X,D_{X})$.
-For example, they have the same orbifold fundamental group, and the same dual complex (up to a PL-homeomorphism) (cf.
-[dFKX17, Proposition 11]).
-On the other hand, the fibration $g:(W,D_{W})\to\mathbb{P}_{A}$ has nice properties, which may reduce some questions about $(X,D_{X})$ to the case $q(X,D_{X})=0$ by induction on dimension.
-We give an example in the last section.
+    $\Delta_{X}=\lfloor D_{X}\rfloor=D_{X}$.
+  2. (2)
+    The quasi-Albanese morphism
+    $\alpha:X\setminus\operatorname{nklt}(X,D_{X})\to(\mathbb{C}^{*})^{n}$
+    is birational and isomorphic in codimension one.
+  3. (3)
+    There is a crepant birational map
+    $(X,D_{X})\dashrightarrow(\mathbb{P}^{1}_{z_{1}}\times\cdots\times\mathbb{P}^{1}_{z_{n}},\sum_{i}((z_{i}=0)+(z_{i}=\infty))).$
+  4. (4)
+    The pair $(X,D_{X})$ is toric.
+A result analogous to Theorem 1.1 or Theorem 1.2 was obtained by Cadorel, Deng, and Yamanoi [Deng2024HyperFund] for Campana’s special manifolds.
+Recall that projective manifolds with nef anticanonical bundle are special (cf.
+[Campana2004Orbifolds, Example 2.3]).
+Moreover, Shokurov’s conjecture [Shokurov2000CompleSurf] only assumes the nefness of $(K_{X}+D_{X})$.
+It is therefore natural to expect that the main Theorems of this note remain valid for log canonical pairs with $(K_{X}+D_{X})$ nef.
+The note is organized as follows. In Section 2, we recall the relevant definitions and develop the necessary technical tools. In Section 3, we prove the main theorems.
 Acknowledgments
-The author thanks professors Zhan Li, Lei Zhang, and Junchao Shentu for useful discussions. The author is supported by USTC.
+The author thanks professors Zhan Li, Lei Zhang, Junchao Shentu, and Osamu Fujino for helpful discussions or comments.
+The author thanks professor Ya Deng for drawing his attention to related results in joint work with Benoit Cadorel and Katsutoshi Yamanoi concerning hyperbolicity and fundamental groups for quasi-projective varieties.
+The author is supported by USTC.
 2. Preparations
-We work over the field of complex numbers. For the singularities of pairs and the minimal model program, we refer to the book [KM98].
-A projective surjective morphism between normal varieties with connected fibers is called a fibration.
-Let $f:X\to Y$ be a fibration between normal varieties.
+We work over the field of complex numbers. For the singularities of pairs and the minimal model program, we refer to [KM98, Kollar2013sing].
+A dominant morphism between normal varieties with irreducible general fibers is called a fibration.
+Let $f:X\to Y$ be a projective fibration between normal varieties.
 Let $D$ be an effective $\mathbb{Q}$-divisor on $X$.
@@
 $D^{v}$ (resp.
-$D^{h}$) is called the vertical part (resp.
+$D^{h}$) is called the vertical (resp.
 horizontal) part of $D$.
@@
 $D=D^{h}$).
-Assume that $D$ is vertical, then $D$ is called exceptional over $Y$ if $\operatorname{codim}f(D)\geq 2$, and very exceptional over $Y$ if for any prime divisor $P\subset Y$, there is a prime divisor $Q\subset X$ such that $Q$ dominates $P$ but $Q\not\subset D$.
-For a birational map $\pi:X\dashrightarrow Y$ between normal varieties, we set $\operatorname{Ex}(\pi):=\sum P$, where the sum is taken over all prime divisors on $X$ which are exceptional over $Y$.
+Assume that $D$ is vertical, then $D$ is called exceptional over $Y$ if $\operatorname{codim}f(D)\geq 2$, and degenerate over $Y$ if for any prime divisor $P\subset Y$, there is a prime divisor $Q\subset X$ such that $Q$ dominates $P$ but $Q\not\subset D$.
+For a birational map $\pi:X\dashrightarrow Y$ between normal varieties, we set $\operatorname{Ex}(\pi)=\sum P$, where the sum is taken over all prime divisors on $X$ which are exceptional over $Y$.
 A birational map $\pi:X\dashrightarrow Y$ is called a birational contraction if $\operatorname{Ex}(\pi^{-1})=0$.
@@
 Let $f:Y\to X$ be a projective surjective morphism from another projective variety, then $f^{-1}\operatorname{Supp}A=\operatorname{Supp}f^{*}A$.
-For a projective variety $X$, we set $q(X):=\dim H^{1}(O_{X})$, and for a projective lc pair $(X,D_{X})$, we set $q(X,D_{X}):=\dim H^{0}(\Omega_{X^{0}}^{1}(\log\lfloor D_{X}|_{X^{0}}\rfloor))$, where $X^{0}$ is the log smooth locus of $(X,D_{X})$.
+For a projective variety $X$, we set $q(X)=\dim H^{1}(X,O_{X})$, and for a projective lc pair $(X,D_{X})$, we set $q(X,D_{X})=\dim H^{0}(X^{0},\Omega_{X^{0}}^{1}(\log\lfloor D_{X}|_{X^{0}}\rfloor))$, where $X^{0}$ is the log smooth locus of $(X,D_{X})$.
 2.1. Differential forms on lc pairs
-Theorem 4 ([GKKP11]).
+Theorem 2.1 ([greb2011differential]).
 Let $(X,D_{X})$ be an lc pair.
-Let $X^{0}$ be its log smooth locus and let $\mathrm{nonklt}(X,D_{X})$ be its non-klt locus.
-Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}$ be the codimension one part of $\pi^{-1}\mathrm{nonklt}(X,D_{X})$.
+Let $X^{0}$ be its log smooth locus and let $\operatorname{nklt}(X,D_{X})$ be its non-klt locus.
+Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}$ be the codimension one part of $\pi^{-1}\operatorname{nklt}(X,D_{X})$.
 Then $\pi_{*}\Omega_{Y}^{k}(\log D_{Y})$ is reflexive and $H^{0}(Y,\Omega_{Y}^{k}(\log D_{Y}))\cong H^{0}(X^{0},\Omega_{X^{0}}^{k}(\log\lfloor D_{X}|_{X^{0}}\rfloor))$ for all $k$.
-Proposition 5.
-Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced. Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Then
+We will use the following Lemma repeatedly.
+Lemma 2.2.
+Let $(X,D_{X})$ be a projective lc pair. Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Then the following holds.
   1. (1)
-    $q(X)=q(Y)$,
+    $q(X,D_{X})=q(Y,D_{Y})$.
   2. (2)
-    $q(X,D_{X})=q(Y,D_{Y})$,
+    $\kappa(K_{X}+D_{X})=\kappa(K_{Y}+D_{Y})$ and $\nu(K_{X}+D_{X})=\nu(K_{Y}+D_{Y})$, where $\kappa$ and $\nu$ are Kodaira dimension and numerical Kodaira dimension respectively.
   3. (3)
-    $\kappa(K_{X}+D)=\kappa(K_{Y}+D_{Y})$ and $\nu(K_{X}+D)=\nu(K_{Y}+D_{Y})$,
-where $\kappa$ and $\nu$ are Kodaira dimension and numerical Kodaira dimension respectively.
+    If $(X,D_{X})$ is dlt, then $q(X)=q(Y)$.
 Proof.
-Since klt varieties have rational singularities, we have (1).
 Since $(X,D_{X})$ is lc, there is an effective exceptional divisor $E$ on $Y$ such that $K_{Y}+D_{Y}=\pi^{*}(K_{X}+D_{X})+E$, so we have (2).
-Now we prove (2). Note that $D_{Y}\geq\pi^{-1}D_{X}$ and $D_{X}=\mathrm{nonklt}(X,D_{X})$. By Theorem 4, to show (2), it suffices to show that the injective homomorphism
-$\iota:H^{0}(\Omega_{Y}^{1}(\log\pi^{-1}D_{X}))\to H^{0}(\Omega_{Y}^{1}(\log D_{Y}))$
-is an isomorphism. Write $\pi^{-1}D_{X}=\sum_{i}P_{i}$ and $D_{Y}-\pi^{-1}D_{X}=\sum_{j}E_{j}$ as sum of prime divisors. By the exact sequence
+If $(X,D_{X})$ is dlt, then $X$ has rational singularities by [KM98, Theorem 5.22].
+Hence we have (3).
+Let $D_{Y}^{\prime}$ be the codimension one part of $\pi^{-1}\operatorname{nklt}(X,D_{X})$.
+Note that $D_{Y}^{\prime}\leq\lfloor D_{Y}\rfloor$.
+To show (1), it suffices to show that the injective homomorphism
+$\iota:H^{0}(Y,\Omega_{Y}^{1}(\log D_{Y}^{\prime}))\to H^{0}(Y,\Omega_{Y}^{1}(\log\lfloor D_{Y}\rfloor))$
+is an isomorphism.
+This isomorphism is a special case of [greb2011differential, Proposition 19.1].
+Here we give this isomorphism a proof using Hacon’s $\mathbb{Q}$-factorial dlt modification [Hacon2014ACC, Proposition 3.3.1].
+Step 1.
+Assume that $(X,D_{X})$ is $\mathbb{Q}$-factorial dlt.
+Then $\operatorname{nklt}(X,D_{X})=\lfloor D_{X}\rfloor$ and $D_{Y}^{\prime}=q^{-1}\lfloor D_{X}\rfloor$.
+Write $D_{Y}^{\prime}=\sum_{i}P_{i}$ and $\lfloor D_{Y}\rfloor-D_{Y}^{\prime}=\sum_{j}E_{j}$ as sums of prime divisors.
+By the exact sequence
 $0\to\Omega_{Y}^{1}\to\Omega_{Y}^{1}(\log(\sum_{i}P_{i}+\sum_{j}E_{j}))\to\oplus_{i}O_{P_{j}}\oplus\oplus_{j}O_{E_{j}}\to 0,$
-one checks that $\iota$ is an isomorphism iff $\sum_{i}a_{i}c_{1}(O_{Y}(P_{i}))+\sum_{j}b_{j}c_{1}(O_{Y}(E_{j}))=0$ implies that all $b_{j}$’s are zero.
-Which is indeed the case since all $E_{j}$’s are exceptional and $X$ is $\mathbb{Q}$-factorial.∎
-Lemma 6.
-Let $(X,D_{X})$ and $(Y,D_{Y})$ be two projective $\mathbb{Q}$-factorial dlt pairs with reduced $D_{X}$ and $D_{Y}$.
+one checks that $\iota$ is an isomorphism iff $\sum_{i}a_{i}c_{1}(O_{Y}(P_{i}))+\sum_{j}b_{j}c_{1}(O_{Y}(E_{j}))=0$ implies that all $b_{j}$’s are zero (see Lemma 3.7).
+This is indeed the case since all $E_{j}$’s are exceptional and $X$ is $\mathbb{Q}$-factorial.
+Step 2.
+Assume that $(X,D_{X})$ is log smooth.
+By Iitaka [iitaka1977logarithmicKodaira, Proposition 1], we have $q(X,\lfloor D_{X}\rfloor)=q(Y,\pi^{-1}\lfloor D_{X}\rfloor)$.
+Hence we have $q(X,\lfloor D_{X}\rfloor)=q(Y,\lfloor D_{Y}\rfloor)$ by Step 1, i.e., (1) holds for log smooth pairs.
+Step 3. In general, consider the following diagram.
+[…]
+where
+  • $(Z,D_{Z})$ is a good minimal model of $(Y,D_{Y})$ over $X$, hence a Hacon’s $\mathbb{Q}$-factorial dlt modification of $(X,D_{X})$ [hacon2007shokurov, Proposition 3.3.1], and
+  • $Y\xleftarrow{p}W\xrightarrow{q}Z$ is a common log resolution.
+Since $K_{Z}+D_{Z}=\mu^{*}(K_{X}+D_{X})$, we have $\mu(\lfloor D_{Z}\rfloor)=\operatorname{nklt}$.
+Hence $\lfloor D_{Z}\rfloor\subset\mu^{-1}\operatorname{nklt}$ and $q^{-1}\lfloor D_{Z}\rfloor\subset(\mu\circ q)^{-1}\operatorname{nklt}$.
+Let $D_{W}^{\prime}$ be the codimension one part of $(\mu\circ q)^{-1}\operatorname{nklt}$, then $q^{-1}\lfloor D_{Z}\rfloor\leq D_{W}^{\prime}$.
+Let $D_{W}=q^{-1}_{*}D_{Z}+\operatorname{Ex}(q)$.
+Then $D_{W}=(\mu\circ q)^{-1}_{*}D_{X}+\operatorname{Ex}(\mu\circ q)$.
+By $q^{-1}\lfloor D_{Z}\rfloor\leq D_{W}^{\prime}\leq\lfloor D_{W}\rfloor$ we have $q(W,q^{-1}\lfloor D_{Z}\rfloor)\leq q(W,D_{W}^{\prime})\leq q(W,\lfloor D_{W}\rfloor)$.
+By Step 1, we have $q(W,q^{-1}\lfloor D_{Z}\rfloor)=q(W,\lfloor D_{W}\rfloor)$.
+Hence
+(2.1.1) $q(W,D_{W}^{\prime})=q(W,\lfloor D_{W}\rfloor).$
+By construction, we have $p^{-1}D_{Y}^{\prime}\subset D_{W}^{\prime}\subset p^{-1}_{*}D_{Y}^{\prime}+\operatorname{Ex}(p)$.
+Hence $q(W,p^{-1}D_{Y}^{\prime})\leq q(W,D_{W}^{\prime})\leq q(W,p^{-1}_{*}D_{Y}^{\prime}+\operatorname{Ex}(p))$.
+By step 1, we have $q(W,p^{-1}D_{Y}^{\prime})=q(W,p^{-1}_{*}D_{Y}^{\prime}+\operatorname{Ex}(p))$.
+Hence $q(W,D_{W}^{\prime})=q(W,p^{-1}D_{Y}^{\prime})$.
+By Iitaka [iitaka1977logarithmicKodaira, Proposition 1], we have
+(2.1.2) $q(W,D_{W}^{\prime})=q(W,p^{-1}D_{Y}^{\prime})=q(Y,D_{Y}^{\prime})$
+By Step 2, we have
+(2.1.3) $q(W,\lfloor D_{W}\rfloor)=q(Y,\lfloor D_{Y}\rfloor)$
+Hence $q(Y,D_{Y}^{\prime})=q(Y,\lfloor D_{Y}\rfloor)$ by the three equations 2.1.1, 2.1.2 and 2.1.3. ∎
+Lemma 2.3.
+Let $(X,D_{X})$ and $(Y,D_{Y})$ be two projective lc pairs.
 Assume that there is a birational contraction $\pi:X\dashrightarrow Y$ such that $D_{X}=\pi^{-1}_{*}D_{Y}+\operatorname{Ex}(\pi)$ then $q(X,D_{X})=q(Y,D_{Y})$.
 Proof.
-Let $X\xleftarrow{p}W\xrightarrow{q}Y$ be a common log resolution.
-Then $q^{-1}_{*}D_{Y}+\operatorname{Ex}(q)=p^{-1}_{*}D_{X}+\operatorname{Ex}(p)$.
-Hence, the Lemma follows from the previous proposition.
-∎
-2.2. Quasi-Albanese morphism
-Theorem 7 (Quasi-Albanese morphism [Iit76], see also [Fuj24]).
+Let $X\xleftarrow{p}W\xrightarrow{q}Y$ be a common log resolution. Then $q^{-1}_{*}D_{Y}+\operatorname{Ex}(q)=p^{-1}_{*}D_{X}+\operatorname{Ex}(p)$. Hence, the Lemma follows from Lemma 2.2. ∎
+2.2. Quasi-Albanese morphisms for smooth varieties
+Theorem 2.4 ([iitaka1976logarithmic], see also [fujino2024quasi]).
 Let $X$ be a smooth quasi-projective variety. Then there is a morphism $\alpha:X\to G$ to a quasi-abelian variety such that
@@
 (2) $\alpha$ is uniquely determined.
-8Construction.
-We recall the constructions of quasi-Albanese morphisms in [Iit76, Fuj24].
+We recall the constructions of quasi-Albanese morphisms in [iitaka1976logarithmic, fujino2024quasi].
 Let $(X,D_{X}=\sum_{i=1}^{m}D_{i})$ be a projective log smooth pair with $D_{X}$ reduced.
 Where $D_{i}$’s are prime divisors.
-Let $X^{0}:=X\setminus D_{X}$ and let $\iota:X^{0}\to X$ be the open embedding.
+Let $X^{0}=X\setminus D_{X}$ and let $\iota:X^{0}\to X$ be the open embedding.
 Let $p:X\to A$ be the Albanese morphism.
 Set $q=q(X)$, $\overline{q}=q(X,D_{X})$, and $d=\overline{q}-q$.
-Let $\eta_{i}$, $1\leq i\leq d$, be a basis of the free part of $\ker(\iota_{*}:H_{1}(X^{0},\mathbb{Z})\to H_{1}(X,\mathbb{Z}))$, and let $\xi_{i}\in H_{1}(X^{0},\mathbb{Z})$, $1\leq i\leq 2q$, be such that these $\eta_{i}$’s and $\xi_{i}$’s form a basis of the free part of $H_{1}(X^{0},\mathbb{Z})$.
-Let $\omega_{i}$, $1\leq i\leq q$, be a basis of $H^{0}(\Omega_{X}^{1})$, and let $\phi_{i}\in H^{0}(\Omega_{X}^{1}(\log D))$, $1\leq i\leq d$, be such that these $\omega_{i}$’s and $\phi_{i}$’s form a basis of $H^{0}(\Omega_{X}^{1}(\log D_{X}))$, and satisfy $\int_{\eta_{i}}\omega_{j}=0$ and $\int_{\eta_{i}}\phi_{j}=\delta_{ij}$ (the Kronecker delta).
+Let $\eta_{i}$, $1\leq i\leq d$, be a basis of the free part of $\ker(\iota_{*}:H_{1}(X^{0},\mathbb{Z})\to H_{1}(X,\mathbb{Z}))$, and let $\xi_{i}\in H_{1}(X^{0},\mathbb{Z})$, $1\leq i\leq 2q$, be such that the $\eta_{i}$ and $\xi_{i}$ form a basis of the free part of $H_{1}(X^{0},\mathbb{Z})$.
+Let $\omega_{i}$, $1\leq i\leq q$, be a basis of $H^{0}(\Omega_{X}^{1})$, and let $\phi_{i}\in H^{0}(\Omega_{X}^{1}(\log D))$, $1\leq i\leq d$, be such that the $\omega_{i}$ and $\phi_{i}$ form a basis of $H^{0}(\Omega_{X}^{1}(\log D_{X}))$, and satisfy $\int_{\eta_{i}}\omega_{j}=0$ and $\int_{\eta_{i}}\phi_{j}=\delta_{ij}$ (the Kronecker delta).
 Set
-$A_{i}:=(\int_{\xi_{i}}\omega_{1},\cdots,\int_{\xi_{i}}\omega_{q},\int_{\xi_{i}}\phi_{1},\cdots,\int_{\xi_{i}}\phi_{d})=:(\hat{A}_{i},C_{i}),1\leq i\leq 2q,\text{ and }$
-$B_{i}:=(\int_{\eta_{i}}\omega_{1},\cdots,\int_{\eta_{i}}\omega_{q},\int_{\eta_{i}}\phi_{1},\cdots,\int_{\eta_{i}}\phi_{d})=:(0,\hat{B}_{i}),1\leq i\leq d.$
+$A_{i}=(\int_{\xi_{i}}\omega_{1},\cdots,\int_{\xi_{i}}\omega_{q},\int_{\xi_{i}}\phi_{1},\cdots,\int_{\xi_{i}}\phi_{d})=:(\hat{A}_{i},C_{i}),1\leq i\leq 2q,\text{ and }$
+$B_{i}=(\int_{\eta_{i}}\omega_{1},\cdots,\int_{\eta_{i}}\omega_{q},\int_{\eta_{i}}\phi_{1},\cdots,\int_{\eta_{i}}\phi_{d})=:(0,\hat{B}_{i}),1\leq i\leq d.$
 The quasi-Albanese variety of $X^{0}$ is given by the quasi-abelian variety
-$G:=\mathbb{C}^{\overline{q}}/\sum_{i}\mathbb{Z}A_{i}+\sum_{j}\mathbb{Z}B_{j}.$
+$G:=\frac{H^{0}(X,\Omega_{X}^{1}(\log D))^{\vee}}{H_{1}(X^{0},\mathbb{Z})}=\frac{\mathbb{C}^{\overline{q}}}{\sum_{i}\mathbb{Z}A_{i}+\sum_{j}\mathbb{Z}B_{j}}.$
 We have the following Chevalley decomposition of $G$:
-$0\to\mathbb{C}^{d}/\sum_{j}\mathbb{Z}\hat{B}_{j}\cong(\mathbb{C}^{*})^{d}\to G\to\mathbb{C}^{q}/\sum_{i}\mathbb{Z}\hat{A}_{i}=A\to 0,$
-and $A$ coincides with the Albanese variety of $X$. The isomorphism $\mathbb{C}^{d}/\sum_{j}\mathbb{Z}\hat{B}_{j}\cong(\mathbb{C}^{*})^{d}$ is given by
+$0\to{\mathbb{C}^{d}}/{\sum_{j}\mathbb{Z}\hat{B}_{j}}\cong(\mathbb{C}^{*})^{d}\to G\to{\mathbb{C}^{q}}/{\sum_{i}\mathbb{Z}\hat{A}_{i}}\to 0,$
+where ${\mathbb{C}^{q}}/{\sum_{i}\mathbb{Z}\hat{A}_{i}}\cong A$, the Albanese variety of $X$, and the isomorphism $\mathbb{C}^{d}/\sum_{j}\mathbb{Z}\hat{B}_{j}\cong(\mathbb{C}^{*})^{d}$ is given by
 $(x_{1},\cdots,x_{d})\mapsto(\exp(2\pi\mathrm{i}x_{1}),\cdots,(\exp(2\pi\mathrm{i}x_{d})).$
@@
 $X^{0}\to G,x\mapsto\left(\int_{x_{0}}^{x}\omega_{1},\cdots,\int_{x_{0}}^{x}\omega_{q},\int_{x_{0}}^{x}\phi_{1},\cdots,\int_{x_{0}}^{x}\phi_{d}\right).$
-Remark 9.
-Keep the notations introduced in the previous construction.
+Since $G$ is quasi-abelian, there are $d=\dim G-\dim A$ line bundles $L_{j}\in\text{Pic}^{0}(A)$ such that $G$ is the fiber product over $A$ of the total spaces of the $L_{j}$ with their zero sections removed (cf.
+[Vojta1996Integral, Section 2]).
+The following content in this subsection aims to explore these line bundles $L_{j}$.
+The explicit form of $L_{j}$, together with Lemma 3.7, helps us calculate examples.
 Let $G^{\prime}\to\mathbb{C}^{q}$ be the pullback of the $(\mathbb{C}^{*})^{d}$-bundle $G\to A$ by the cover $\mathbb{C}^{q}\to A$.
@@
 $\pi_{1}(A)\to\operatorname{Aut}((\mathbb{C}^{*})^{d}),\hat{A}_{i}\mapsto\left((t_{1},\cdots,t_{d})\mapsto(t_{1}\exp(2\pi\mathrm{i}\int_{\xi_{i}}\phi_{1}),\cdots,t_{d}\exp(2\pi\mathrm{i}\int_{\xi_{i}}\phi_{d}))\right),1\leq i\leq q.$
-For each $j$, let $L_{j}\in\mathrm{Pic}^{0}(A)$ be the line bundle corresponding to the representation
-(1) $\pi_{1}(A)\to\mathbb{C}^{*},\hat{A}_{i}\mapsto\exp(2\pi\mathrm{i}\int_{\xi_{i}}\phi_{j}),1\leq i\leq q.$
-Then $G$ is the fiber product over $A$ of the total spaces of $L_{j}$’s with zero sections removed, in other words,
-$G=\operatorname{Spec}_{A}\bigoplus_{(m_{1},\cdots,m_{d})\in\mathbb{Z}^{d}}L_{1}^{-m_{1}}\otimes\cdots\otimes L_{d}^{-m_{d}}.$
-$G$ has the following natural compactification:
-(2) $G\subset\mathbb{P}_{A}:=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1}).$
+Hence $L_{j}$ is the line bundle corresponding to the representation
+(2.2.1) $\pi_{1}(A)\to\mathbb{C}^{*},\hat{A}_{i}\mapsto\exp(2\pi\mathrm{i}\int_{\xi_{i}}\phi_{j}),1\leq i\leq q.$
 Let $a_{ij}=2\pi\mathrm{i}\mathrm{Res}_{D_{i}}\phi_{j}$, $1\leq i\leq m$, $1\leq j\leq d$. For each $j$, we have the following.
@@
 Since $\pi_{1}(X^{0})\to\pi_{1}(X)$ is surjective with kernel generated by $\delta_{i}$’s, $\exp(2\pi\mathrm{i}\int_{\delta_{i}}\phi_{j})=1$ for all $i$, and $\phi_{j}$ is $d$-closed, the expression $\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})$ is well-defined.
-Hence, we have (ii) by the representation (1).
-(3) By $t_{j}(\gamma x)=\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})t_{j}(x)$ for all $\gamma\in\pi_{1}(X)$, we have $t_{j}$ is a holomorphic nowhere vanishing section of $p^{*}L_{j}$ over $X^{0}$.
-Let $\{x_{i}\}$ be a coordinate system on an open subset $U$ of $X$ such that $D=(x_{1}x_{2}\cdots x_{c}=0)$.
+Hence, we have (2) by the representation (2.2.1).
+(3) Since $t_{j}(\gamma x)=\exp(2\pi\mathrm{i}\int_{\gamma}\phi_{j})t_{j}(x)$ for all $\gamma\in\pi_{1}(X)$, we have $t_{j}$ is a holomorphic nowhere vanishing section of $p^{*}L_{j}$ over $X^{0}$.
+Let $\{x_{i}\}$ be a coordinate system on an analytic open subset $U$ of $X$ such that $D=(x_{1}x_{2}\cdots x_{c}=0)$.
 Then
@@
 $t_{j}|_{U}=\prod_{i=1}^{c}x_{i}^{a_{ij}}\times\text{some holomorphic function without zeros}.$
-Hence $t_{j}$ is a meromorphic, and so rational (by Serre’s GAGA) section of $p^{*}L_{j}$ and $\mathrm{div}(t_{j})=\sum_{i}a_{ij}D_{i}$. ∎
+Hence $t_{j}$ is a meromorphic, and so a rational (by Serre’s GAGA) section of $p^{*}L_{j}$ and $\mathrm{div}(t_{j})=\sum_{i}a_{ij}D_{i}$. ∎
+Let $(\mathbb{P}_{A},H)$ be the canonical compactification of $G$.
 Let $(z_{1},\cdots,z_{d})$ be the inhomogeneous fiber coordinates of the $(\mathbb{P}^{1})^{d}$-bundle $\mathbb{P}_{A}\to A$, the rational map $X\dashrightarrow\mathbb{P}_{A}$ over $A$ is now given by
 $x\mapsto(z_{1},\cdots,z_{d})=(t_{1},\cdots,t_{d})\in\mathbb{P}_{A}.$
-Let $H:=\mathbb{P}_{A}\setminus G=\sum_{i}((z_{i}=0)+(z_{i}=\infty))$. The pair $(\mathbb{P}_{A},H)$ is log smooth and $K_{\mathbb{P}_{A}}+H\sim 0$.
-Example 10.
+Example 2.5.
 Let $(X,D_{X})$ be a log smooth pair with $D_{X}$ reduced.
 (1) Let $X=\mathbb{P}^{2}$.
-If $D_{X}=$ 3 lines, then $q(X,D_{X})=3$.
+If $D_{X}=$ 3 lines, then $q(X,D_{X})=3$, and $X\setminus D_{X}=(\mathbb{C}^{*})^{2}$ itself is a quasi-abelian variety.
 If $D_{X}=$ an elliptic curve, then $q(X,D_{X})=0$.
 If $D_{X}=Q+L$, where $L=(l=0)$ is a line and $Q=(q=0)$ is a conic, then $q(X,D_{X})=1$, $H^{0}(\Omega^{1}_{X}(\log D_{X}))=\mathbb{C}\cdot d\log\frac{l^{2}}{q}$ and the quasi-Albanese morphism is $X\setminus D\to\mathbb{C}^{*},x\mapsto\frac{l^{2}}{q}$.
+The rational map $X\dashrightarrow\mathbb{P}^{1}$, $x\mapsto[l^{2}:q]$ is not defined at $Q\cap L$.
 (2) Let $X=\mathbb{P}^{1}_{z_{1}}\times\mathbb{P}^{1}_{z_{2}}$ and $D_{X}=F_{1}+F_{2}+\Delta$ where $F_{1}=(z_{1}=0)$, $F_{2}=(z_{2}=0)$, $\Delta=(z_{1}=z_{2})$.
-Then $H^{0}(\Omega_{X}(\log D_{X}))=\mathbb{C}\cdot d\log(1/z_{1}-1/z_{2})$ and the quasi-Albanese morphism is $X\setminus D\to\mathbb{C}^{*},(z_{1},z_{2})\mapsto 1/z_{1}-1/z_{2}$.
+Then $H^{0}(\Omega_{X}^{1}(\log D_{X}))=\mathbb{C}\cdot d\log(1/z_{1}-1/z_{2})$ and the quasi-Albanese morphism is $X\setminus D\to\mathbb{C}^{*},(z_{1},z_{2})\mapsto 1/z_{1}-1/z_{2}$.
+The rational map $X\dashrightarrow\mathbb{P}^{1}$, $(z_{1},z_{2})\mapsto 1/z_{1}-1/z_{2}$ is not defined at $(F_{1}+F_{2})\cap\Delta=(z_{1}=z_{2}=0)$.
 (3) Let $A$ be an abelian variety and let $L$ be a line bundle on $A$.
@@
 If $L\not\in\text{Pic}^{0}(A)$, then $q(X,D_{X})=\dim A$, and the quasi-Albanese morphism is the natural projection $\mathbb{P}_{A}\setminus H_{1}+H_{2}\to A$.
-Remark 11.
+Lemma 2.6.
+Let $X$ be a smooth projective variety. Let $D_{X}\leq D_{X}^{\prime}$ be reduced snc divisors. By the universal property for quasi-Albanese morphisms, we have the following commutative diagram.
+$X\setminus D_{X}^{\prime}$$G^{\prime}$$X\setminus D_{X}$}G{G}
+where the horizontal arrows are quasi-Albanese morphisms and $G^{\prime}\to G$ is the induced morphism. Assume that $q(X,D_{X})=q(X,D_{X}^{\prime})$. Then $G^{\prime}\to G$ is an isomorphism.
+Proof.
+Since the injective homomorphism
+$H^{0}(X,\Omega_{X}^{1}(\log D_{X}))\to H^{0}(X,\Omega_{X}^{1}(\log D_{X}^{\prime}))$
+is an isomorphism, we have $H^{1}(X\setminus D_{X},\mathbb{C})\to H^{1}(X\setminus D_{X}^{\prime},\mathbb{C})$ is an isomorphism by Deligne’s mixed Hodge structures:
+$H^{1}(X\setminus D_{X},\mathbb{C})$$H^{1}(X\setminus D_{X}^{\prime},\mathbb{C})$$H^{1}(X,O_{X})\oplus H^{0}(X,\Omega_{X}^{1}(\log D_{X}))$$H^{1}(X,O_{X})\oplus H^{0}(X,\Omega_{X}^{1}(\log D_{X}^{\prime}))$$\scriptstyle{\cong}\scriptstyle{\cong}\scriptstyle{\cong}$
+Hence
+$G^{\prime}=\frac{H^{0}(X,\Omega_{X}^{1}(\log D_{X}^{\prime}))^{\vee}}{H_{1}(Z\setminus D_{X}^{\prime},\mathbb{Z})}\to G=\frac{H^{0}(X,\Omega_{X}^{1}(\log D_{X}))^{\vee}}{H_{1}(Z\setminus D_{X},\mathbb{Z})}$
+is an isomorphism. ∎
+2.3. Quasi-Albanese morphisms for klt varieties
 Let $X$ be a (possibly) singular variety.
@@
 Assume that $X$ has klt singularities.
-Then the fibers of $\pi$ are rationally chain connected by Hacon-McKernan [HM07].
-Since $G$ contains no rational curves, $f^{\prime}$ factors through $\pi$ by the rigidity Lemma (cf.[Deb01, Lemma 1.15]).
-The statement in [Deb01, Lemma 1.15] requires $f^{\prime}$ to be proper, but the proof only used that the image of $X^{\prime}\to X\times G$ is proper over $X$, which is guaranteed by [Har77, Chapter II Exercise 4.4].
-Proposition 12.
-Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $\kappa(K_{X}+D_{X})=0$.
-Then the quasi-Albanese morphism $X\setminus D_{X}\to G$ is dominant and has irreducible general fibers.
+Then the fibers of $\pi$ are rationally chain connected by Hacon-McKernan [hacon2007shokurov].
+Since $G$ contains no rational curves, $f^{\prime}$ factors through $\pi$ by the rigidity Lemma (cf.
+[KollarRationalCurves, 5.3 Proposition]).
+In other words, there is a morphism $f:X\to G$ such that $f\circ\pi=f^{\prime}$.
+[…]
+One checks that the morphism $f:X\to G$ given in this way satisfies the universal property as in Theorem 2.4.
+Let $(X,D_{X})$ be a projective lc pair and let $\operatorname{nklt}=\operatorname{nklt}(X,D_{X})$ be its non-klt locus.
+Let $\alpha:X\setminus\operatorname{nklt}\to G$ be the quasi-Albanese morphism.
+Let $A$ be the compact part of $G$ and let $A^{\prime}:=(\text{Pic}^{0}(X))^{\vee}_{\text{red}}$ be the Albanese variety of $X$.
+Let $\pi:Z\to X$ be a log resolution such that $\pi^{-1}\operatorname{nklt}$ is a divisor.
+By our definition, $G$ is the quasi-Albanese variety of $Z\setminus\pi^{-1}\operatorname{nklt}$.
+Hence $A$ is the Albanese variety of $Z$ by Iitaka’s construction.
+By the universal property for Albanese morphisms, we have the following commutative diagram.
+[…]
+where $A\to A^{\prime}$ is the morphism induced by $Z\to X$.
+Unlike the log smooth case, in general, we have $A\not\cong A^{\prime}$.
+But if $(X,D_{X})$ is dlt, then by [hacon2007shokurov] and the rigidity Lemma, we do have $A\cong A^{\prime}$.
+Lemma 2.7.
+Let $(X,D_{X})$ be a projective lc pair.
+Let $G$ be the quasi-Albanese variety of $X\setminus\operatorname{nklt}(X,D_{X})$.
+Let $\pi:Z\to X$ be a log resolution such that $D_{Z}^{\prime}:=\pi^{-1}\operatorname{nklt}(X,D_{X})$ is a divisor.
+Let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$.
+Note that $D_{Z}^{\prime}\leq\lfloor D_{Z}\rfloor$.
+By the universal property for quasi-Albanese morphisms, we have the following commutative diagram.
+$Z\setminus\lfloor D_{Z}\rfloor$$G^{\prime}$$Z\setminus D_{Z}^{\prime}$}G{G}
+where the horizontal arrows are quasi-Albanese morphisms and $G^{\prime}\to G$ is the induced morphism. Then $G^{\prime}\to G$ is an isomorphism. In particular, we have $\dim G=q(X,D_{X})$.
 Proof.
-Let $\pi:Y\to X$ be a log resolution of $(X,D_{X})$ and let $D_{Y}:=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$.
-Then $q(Y,D_{Y})=q(Y,\pi^{-1}D_{X})=q(X,D_{X})$ and $\kappa(K_{Y}+D_{Y})=0$ by Proposition 5.
-By the universal property of quasi-Albanese morphisms, we have the following commutative diagram.
+By Lemma 2.2, we have $q(X,D_{X})=q(Z,D_{Z}^{\prime})=q(Z,D_{Z})$. Hence the Lemma follows from Lemma 2.6. ∎
+2.4. Subadditivity and positivity
+The following is a summary of slightly generalizations of the relevant theorems due to Kawamata [kawamata1981characterization], Maehara [maehara1986weak], Campana [Campana2004Orbifolds] and Fujino [Fujino2017notes], and it is one of the most important techniques used in this note.
+Theorem 2.8.
+Let $f:X\to Y$ be a fibration between smooth projective varieties.
+Let $D_{X}$ and $D_{Y}$ be divisors on $X$ and $Y$ respectively, such that $(X,D_{X})$ and $(Y,D_{Y})$ are log smooth lc.
+Assume that $D_{Y}$ is reduced.
+Denote by $F$ a very general fiber of $f$.
+Assume that
+$f^{-1}D_{Y}\subset\lfloor D_{X}\rfloor.$
+Then the following holds.
+(1)(cf.
+[kawamata1981characterization, Theorem 30], [maehara1986weak, Corollary 1, Corollary 2], [Campana2004Orbifolds, Theorem 4.2] or [Fujino2017notes, Theorem 1.9]) Let $B$ be a big $\mathbb{Q}$-divisor on $Y$, then
+$\kappa(K_{X}+D_{X}-f^{*}(K_{Y}+D_{Y})+f^{*}B)=\kappa(K_{F}+{D_{X}}|_{F})+\dim Y.$
+In particular, if $K_{Y}+D_{Y}$ is big, then
+$\kappa(K_{X}+D_{X})=\kappa(K_{F}+D_{X}|_{F})+\dim Y.$
+(2) If $\kappa(K_{F}+{D_{X}}|_{F})\geq 0$, then $K_{X}+D_{X}-f^{*}(K_{Y}+D_{Y})$ is pseudo-effective.
 […]
-where $g$, $g^{\prime}$ and $f$ are quasi-Albanese morphisms.
-Since the injective homomorphism $H^{0}(\Omega^{1}_{Y}(\log\pi^{-1}D_{X}))\to H^{0}(\Omega_{Y}^{1}(\log D_{Y}))$ is an isomorphism, we have $H^{1}(Y\setminus D_{Y},\mathbb{C})=H^{1}(O_{Y})\oplus H^{0}(\Omega_{Y}(\log\pi^{-1}D_{X}))\to H^{1}(Y\setminus\pi^{-1}D_{X},\mathbb{C})=H^{1}(O_{Y})\oplus H^{0}(\Omega_{Y}^{1}(\log D_{Y}))$ is an isomorphism by Deligne’s theory of mixed Hodge structures.
-Hence the surjective morphism $H_{1}(Y\setminus D_{Y},\mathbb{Z})/{\mathrm{torsion}}\to H_{1}(Y\setminus\pi^{-1}D_{X},\mathbb{Z})/{\mathrm{torsion}}$ is also an isomorphism, so $G^{\prime}\to G$ is an isomorphism.
-Since $\kappa(K_{Y}+D_{Y})=0$, we have $g^{\prime}$ is dominant and has irreducible general fibers by Theorem 1. Hence, $g$, $f$ are dominant and have irreducible general fibers. ∎
-2.3. Subadditivity and positivity
-Theorem 13.
-Let $f:X\to Y$ be a fibration between smooth projective varieties.
-Let $D_{X}$ and $D_{Y}$ be reduced snc divisors on $X$ and $Y$ respectively.
-Denote by $F$ a general fiber of $f$.
-Assume that $f^{-1}D_{Y}\subset D_{X}$.
+Proof.
+Note that we do not require $D_{X}$ to be reduced here.
+Using [Fujino2017notes, Theorem 1.1], the proof of (1) is identical to that of [Fujino2017notes, Theorem 1.9].
+Since the $\mathbb{Q}$-divisor $B$ in (1) can be sufficiently small, (2) follows.
+∎
+Let $(X^{0},D^{0})$ be a quasi-projective log smooth lc pair, and let $X^{0}\subset X$ be a compactification such that $(X,D+B)$ is log smooth, where $B=X\setminus X^{0}$ and $D$ is the closure of $D^{0}$ in $X$.
+We define the logarithmic Kodaira dimension $\overline{\kappa}$ (resp.
+numerical logarithmic Kodaira dimension $\overline{\nu}$) of $(X^{0},D^{0})$ as $\overline{\kappa}(X^{0},D^{0}):=\kappa(K_{X}+D+B)$ (resp.
+$\overline{\nu}(X^{0},D^{0}):=\nu(K_{X}+D+B)$).
+To show that $\overline{\kappa}$ and $\overline{\nu}$ are well defined, it suffices to show that any two such compactifications give the same $\overline{\kappa}$ and $\overline{\nu}$.
+Let $(X,D+B)$ and $(X^{\prime},D^{\prime}+B^{\prime})$ be two such compactifications.
+Let $X\xleftarrow{p}X^{\prime\prime}\xrightarrow{q}X^{\prime}$ be a common resolution, which is an isomorphism over $X^{0}$, and such that $D^{\prime\prime}+B^{\prime\prime}$ is snc, where $B^{\prime\prime}=X^{\prime\prime}\setminus X^{0}$ and $D^{\prime\prime}$ is the closure of $D^{0}$ in $X^{\prime\prime}$.
+Then $D^{\prime\prime}=p^{-1}_{*}D$ and $B^{\prime\prime}=p^{-1}_{*}B+\operatorname{Ex}(p)$.
+Hence
+$K_{X^{\prime\prime}}+D^{\prime\prime}+B^{\prime\prime}=p^{*}(K_{X}+D+B)+\sum_{E\subset\operatorname{Ex}(p)}(a(E,X,D+B)+1)E.$
+Since $(X,D+B)$ is lc, we have $\sum_{E\subset\operatorname{Ex}(p)}(a(E,X,D+B)+1)E\geq 0$.
+Hence $\kappa(K_{X}+D+B)=\kappa(K_{X^{\prime\prime}}+D^{\prime\prime}+B^{\prime\prime})=\kappa(K_{X^{\prime}}+D^{\prime}+B^{\prime})$ and $\nu(K_{X}+D+B)=\nu(K_{X^{\prime\prime}}+D^{\prime\prime}+B^{\prime\prime})=\nu(K_{X^{\prime}}+D^{\prime}+B^{\prime})$.
+Using these definitions, we can restate Theorem 2.8 (1) (3) as follows.
+Lemma 2.9.
+Let $f^{0}:X^{0}\to Y^{0}$ be a fibration between smooth quasi-projective varieties.
+Let $D^{0}$ be a divisor on $X^{0}$ such that $(X^{0},D^{0})$ is log smooth lc.
+Denote by $F^{0}$ a very general fiber of $f^{0}$.
 Then the following holds.
-(1)[Mae86, Corollary 2] Let $B$ be a big $\mathbb{Q}$-divisor on $Y$, then
-$\kappa(K_{X}+D_{X}-f^{*}(K_{Y}+D_{Y})+f^{*}B)\geq\kappa(K_{F}+{D_{X}}|_{F})+\dim Y.$
-(2) If $\kappa(K_{F}+{D_{X}}|_{F})\geq 0$, then $K_{X}+D_{X}-f^{*}(K_{Y}+D_{Y})$ is pseudoeffective.
-[…]
-Since the $\mathbb{Q}$-divisor $B$ in (1) can be sufficiently small, (2) follows.
-Let $X^{0}$ be a smooth quasi-projective variety and let $X^{0}\subset X$ be a compactification such that $D_{X}:=X\setminus X^{0}$ is a simple normal crossing divisor.
-By the same reasoning as in the definition of $\overline{\kappa}(X^{0})$ (cf.
-[Fuj24, Lemma 2.3]), we can define the numerical logarithmic Kodaira dimension as $\overline{\nu}(X^{0}):=\nu(K_{X}+D_{X})$.
-By Fujino’s subadditivity 13(3), we have the following.
-Lemma 14.
-Let $f:X\to Y$ be a dominant morphism with irreducible general fibers between smooth quasi-projective varieties.
-Let $F$ be a general fiber of $f$.
-Then $\overline{\nu}(X)\geq\overline{\nu}(F)+\overline{\kappa}(Y)$.
+  1. (1)
+    If $\overline{\kappa}(Y^{0})=\dim Y^{0}$, then $\overline{\kappa}(X^{0},D^{0})=\overline{\kappa}(F^{0},D^{0}|_{F^{0}})+\dim Y^{0}$.
+  2. (2)
+    $\overline{\nu}(X^{0},D^{0})\geq\overline{\nu}(F^{0},D^{0}|_{F^{0}})+\overline{\kappa}(Y^{0})$.
+Proof.
+Let $X^{0}\subset X$ and $Y^{0}\subset Y$ be compactifications, such that $(X,D+B_{X})$, $(Y,B_{Y})$ are log smooth and there is a morphism $f:X\to Y$ extending $f^{0}$, where $D$ is the closure of $D^{0}$ in $X$, $B_{X}=X\setminus X^{0}$ and $B_{Y}=Y\setminus Y^{0}$.
+By construction, we have $f^{-1}B_{Y}\subset B_{X}\subset\lfloor D+B_{X}\rfloor$.
+Then (1), (2) follow from Theorem 2.8 (1), (3) respectively.
+∎
 3. Proof of the main theorems
-Proof of Theorem 2.
-Let $\alpha:X\setminus D_{X}\to G$ be the quasi-Albanese morphism.
-Let $G\subset\mathbb{P}_{A}=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1})$ be the compactification described in remark 9, where $L_{i}$’s are elements in $\text{Pic}^{0}(A)$ and let $p:\mathbb{P}_{A}\to A$ be the natural projection.
-Then $H:=\mathbb{P}_{A}\setminus G$ is a snc divisor and $K_{\mathbb{P}_{A}}+H\sim 0$.
-Let $\pi:Z\to X$ be a log resolution of $(X,D_{X})$ resolving the indeterminacies of the rational map $X\dashrightarrow\mathbb{P}_{A}$ and let $D_{Z}:=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$.
-By construction, we have
-(3) $f^{-1}H\subset D_{Z}.$
-Write $K_{Z}+D_{Z}=\pi^{*}(K_{X}+D_{X})+E_{Z}$.
-Since $(X,D_{X})$ is lc, $E_{Z}$ is effective and exceptional over $X$.
-Then by Theorem 15, $(Z,D_{Z})$ has a good minimal model $(W,D_{W})$ over $\mathbb{P}_{A}$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$, the canonical bundle formula for $(W,D_{W})\to\mathbb{P}_{A}$ writes $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H)$, the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, and $g^{-1}H=D_{W}^{v}$.
-[…]
-Since the birational map $Z\dashrightarrow W$ contracts precisely the divisor $E_{Z}\subset\mathrm{Ex}(\pi)$, we have $\mu:W\dashrightarrow X$ is a birational contraction and $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$.
-Let $X\xleftarrow{u}V\xrightarrow{v}W$ be a common log resolution of $(X,D_{X})$ and $(W,D_{W})$.
-Then $u^{*}(K_{X}+D_{X})=v^{*}(K_{W}+D_{W})$ by the Negativity Lemma, i.e., $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ is crepant.
-By Lemma 6, we have $q(W,D_{W})=q(X,D_{X})$ and $q(W)=q(X)$.
-By the universal property of quasi-Albanese morphisms and Proposition 12, the composite morphism $W\setminus D_{W}\to W\setminus g^{-1}H\to G$ is the quasi-Albanese morphism of $W\setminus D_{W}$.
-By the same reasoning, $p\circ g:W\to A$ is the Albanese morphism of $W$.
+3.1. Proof of theorem 1.2
+Proof of theorem 1.2.
+This is (3) of the following Theorem 3.1. ∎
+Theorem 3.1.
+  1. (1)
+    Let $(X^{0},D_{X}^{0})$ be a quasi-projective log smooth lc pair with $\overline{\kappa}(X^{0},D_{X}^{0})=0$. Then the quasi-Albanese morphism of $X^{0}$ is a fibration.
+  2. (2)
+    Let $(X,D_{X})$ be a projective $\mathbb{Q}$-factorial dlt pair with $\kappa(K_{X}+D_{X})=0$.
+Let $0\leq\Delta_{X}\leq\lfloor D_{X}\rfloor$ be a reduced divisor.
+Then the quasi-Albanese morphism of $X\setminus\Delta_{X}$ is a fibration.
+  3. (3)
+    Let $(X,D_{X})$ be a projective lc pair with $\kappa(K_{X}+D_{X})=0$. Then the quasi-Albanese morphism of $X\setminus\operatorname{nklt}(X,D_{X})$ is a fibration.
+Proof.
+(1) The proof is identical to that of [fujino2024quasi, Theorem 1.2] by using Lemma 2.9 (1) instead of [fujino2024quasi, Theorem 6.1].
+(2) Let $\pi:Z\to X$ be a log resolution and let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$.
+Since $\pi^{-1}\Delta_{X}\leq\lfloor D_{Z}\rfloor$, we have $\overline{\kappa}(Z\setminus\pi^{-1}\Delta_{X},D_{Z}-\pi^{-1}\Delta_{X})=\kappa(K_{Z}+D_{Z})=\kappa(K_{X}+D_{X})=0$.
+By (1), the quasi-Albanese morphism of $Z\setminus\pi^{-1}\Delta_{X}$ is a fibration.
+Hence we have (2).
+(3) Let $\pi:Z\to X$ be a log resolution such that $\pi^{-1}\operatorname{nklt}(X,D_{X})$ is a divisor.
+Let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$.
+Since $\pi^{-1}\operatorname{nklt}\leq\lfloor D_{Z}\rfloor$, we have $\overline{\kappa}(Z\setminus\pi^{-1}\operatorname{nklt},D_{Z}-\pi^{-1}\operatorname{nklt})=\kappa(K_{Z}+D_{Z})=\kappa(K_{X}+D_{X})=0$.
+By (1), the quasi-Albanese morphism of $Z\setminus\pi^{-1}\operatorname{nklt}$ is a fibration.
+Hence we have (3).
 ∎
-Theorem 15.
-Let $(Z,D_{Z})$ be a projective log smooth pair with $D_{Z}$ reduced and $\kappa(K_{Z}+D_{Z})=0$.
-Let $A$ be the Albanese variety of $Z$.
-Let $G$ be a quasi-abelian variety whose abelian part is $A$.
-Let
-$\mathbb{P}_{A}=\mathbb{P}_{A}(O_{A}\oplus L_{1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d})$
-be a compactification of $G$ where $L_{i}$’s are elements in $\text{Pic}^{0}(A)$.
-Let $(z_{1},\cdots,z_{d})$ be the inhomogeneous fiber coordinates of $\mathbb{P}_{A}$.
-For each $i$, let $p_{i}:\mathbb{P}_{A}\to\mathbb{P}_{A}(O_{A}\oplus L_{i})$ be the projection, and let
-$H_{i}=(z_{i}=0)+(z_{i}=\infty)\subset\mathbb{P}_{A}(O_{A}\oplus L_{d})$
-be the effective and ample over $A$ divisor. Let $H=\sum_{i}p_{i}^{*}H_{i}$.
+3.2. Proof of theorem 1.4
+In the following theorem, we do not assume that $G$ is the quasi-Albanese variety of $Z\setminus\lfloor D_{Z}\rfloor$ and we do not assume that $G\neq A$ (in particular, $(\mathbb{P}_{A},H)$ may coincides with $(A,0)$).
+Theorem 3.2.
+Let $(Z,D_{Z})$ be a projective log smooth lc pair with $\kappa(K_{Z}+D_{Z})=0$.
+Let $G$ be a quasi-abelian variety and let $A$ be its compact part.
+Let $(\mathbb{P}_{A},H)$ be the canonical compactification of $G$.
+Note that $(\mathbb{P}_{A},H)$ is log smooth and $K_{\mathbb{P}_{A}}+H\sim 0$.
 Assume that there are
-  • a fibration $f:Z\to\mathbb{P}_{A}$ with $f^{-1}H\subset D_{Z}$,
-  • a projective $\mathbb{Q}$-factorial dlt pair $(X,D_{X})$ with $K_{X}+D_{X}\sim_{\mathbb{Q}}0$, and
-  • a birational morphism $\pi:Z\to X$ such that $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$ and
+  • a fibration $f:Z\to\mathbb{P}_{A}$ with $f^{-1}H\subset\lfloor D_{Z}\rfloor$,
+  • a projective lc pair $(X,D_{X})$ with $K_{X}+D_{X}\sim_{\mathbb{Q}}0$, and
+  • a birational morphism $\pi:Z\to X$ such that $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Note that since $(X,D_{X})$ is lc, we have
     $K_{Z}+D_{Z}=\pi^{*}(K_{X}+D_{X})+E_{Z}$
@@
   1. (1)
-    Any exceptional over $\mathbb{P}_{A}$ divisor is contained in $E_{Z}+f^{-1}H$.
+    Let $Q\subset Z$ be a prime divisor which is exceptional over $\mathbb{P}_{A}$. Then $Q\subset E_{Z}+f^{-1}H$.
   2. (2)
@@
     $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H),$
-    the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, and the vertical part $D_{W}^{v}$ of $D_{W}$ coincides with $g^{-1}H$.
+    the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, and $D_{W}^{v}=\lfloor D_{W}^{v}\rfloor=g^{-1}H$.
+The induced birational map $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ is a crepant birational contraction with $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$.
   5. (5)
@@
 Proof.
-Note that by assumption, $G=\mathbb{P}_{A}\setminus H$, $(\mathbb{P}_{A},H)$ is log smooth, and $K_{\mathbb{P}_{A}}+H\sim 0$.
-Let $F$ be a general fiber of $f$.
-By Fujino’s subadditivity Theorem 13(3), we have
-$0=\nu(K_{Z}+D_{Z})\geq\nu(K_{F}+D_{Z}|_{F})+\kappa(K_{\mathbb{P}_{A}}+H),$
-so $\nu(K_{F}+D_{Z}|_{F})=0$. By Gongyo’s abundance theorem [Gon11], $(F,D_{Z}|_{F})$ has a good minimal model and $\kappa(K_{F}+D_{Z}|_{F})=0$.
-(1) Let $Q\subset Z$ be an exceptional over $\mathbb{P}_{A}$ prime divisor which is not contained in $f^{-1}H$.
+Let $F$ be a very general fiber of $f$. By Fujino’s subadditivity Theorem 2.8(3), we have
+$0=\nu(K_{Z}+D_{Z})\geq\nu(K_{F}+D_{Z}|_{F})+\kappa(K_{\mathbb{P}_{A}}+H).$
+Hence $\nu(K_{F}+D_{Z}|_{F})=0$. By Gongyo’s abundance theorem [gongyo2011minimal], we have $\kappa(K_{F}+D_{Z}|_{F})=0$ and $(F,D_{Z}|_{F})$ has a good minimal model.
+(1) Assume that $Q\not\subset f^{-1}H$.
 Let $q:\mathbb{P}^{\prime}_{A}\to\mathbb{P}_{A}$ be a composition of the blowup along $f(Q)$ with a log resolution.
@@
 Write $K_{Z^{\prime}}+D_{Z^{\prime}}=p^{*}(K_{Z}+D_{Z})+E_{1}$, $K_{\mathbb{P}^{\prime}_{A}}+H^{\prime}=q^{*}(K_{\mathbb{P}_{A}}+H)+E_{2}$, where $H^{\prime}=q^{-1}H$, $D_{Z^{\prime}}=p^{-1}_{*}D_{Z}+\operatorname{Ex}(p)$, and $E_{1}$, $E_{2}$ are effective and exceptional divisors.
-By $f^{\prime-1}H^{\prime}\subset D_{Z^{\prime}}$,
+By $f^{-1}H\subset\lfloor D_{Z}\rfloor$, we have $f^{\prime-1}H^{\prime}\subset\lfloor D_{Z^{\prime}}\rfloor$.
+By
 $0\sim_{\mathbb{Q}}p^{*}(K_{Z}+D_{Z}-f^{*}(K_{\mathbb{P}_{A}}+H)-E_{Z})=K_{Z^{\prime}}+D_{Z^{\prime}}-f^{\prime*}(K_{\mathbb{P}^{\prime}_{A}}+H^{\prime})-E_{1}+f^{\prime*}E_{2}-p^{*}E_{Z},$
-and Theorem 13(1), we have $E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z}$ is pseudoeffective.
-Since $(\pi\circ p)_{*}(E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z})=-(\pi\circ p)_{*}f^{\prime*}E_{2}$ is pseudoeffective, we have $f^{\prime*}E_{2}$ is exceptional over $X$.
+$\kappa(K_{F}+D_{Z}|_{F})=0$, and Theorem 2.8(2), we have $E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z}$ is pseudo-effective.
+Since $(\pi\circ p)_{*}(E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z})=-(\pi\circ p)_{*}f^{\prime*}E_{2}$ is pseudo-effective, we have $f^{\prime*}E_{2}$ is exceptional over $X$.
 Hence $E_{1}-f^{\prime*}E_{2}+p^{*}E_{Z}$ is effective by a Lemma due to Lazarsfeld (cf.
-[KL09, Corollary 13]).
+[kollar2010quotients, Corollary 13]).
 Let $Q^{\prime}$ be the strict transform of $Q$ on $Z^{\prime}$ and let $P^{\prime}=f^{\prime}(Q^{\prime})$.
@@
 for some effective exceptional over $\mathbb{P}_{A}$ divisor $E$. Then by (1), we have
-$f^{-1}(P+H)\subset E_{Z}+f^{-1}H\subset D_{Z}.$
-Then by Lemma 14 and Lemma 16, we have
-$0=\nu(K_{Z}+D_{Z})=\overline{\nu}(Z\setminus D_{Z})\geq\overline{\nu}(F\setminus D_{Z})+\overline{\kappa}({\mathbb{P}_{A}}\setminus P+H)=\overline{\kappa}(G\setminus P)>0,$
+$f^{-1}(P+H)\subset E_{Z}+f^{-1}H\subset\lfloor D_{Z}\rfloor.$
+Consider the fibration $Z\setminus\lfloor D_{Z}\rfloor\to\mathbb{P}_{A}\setminus P+H$. By Lemma 2.9(2) and Lemma 3.2, we have
+$0=\nu(K_{Z}+D_{Z})=\overline{\nu}(Z\setminus\lfloor D_{Z}\rfloor,D_{Z})\geq\overline{\nu}(F\setminus\lfloor D_{Z}\rfloor,D_{Z})+\overline{\kappa}({\mathbb{P}_{A}}\setminus P+H)=\overline{\kappa}(G\setminus P)>0,$
 which is a contradiction.
-(3) By (2) and the assumption in (3), we known that the vertical part $E_{Z}^{v}$ of $E_{Z}$ is very exceptional over $\mathbb{P}_{A}$. Since $\kappa(E_{Z}|_{F})=\kappa(K_{F}+D_{Z}|_{F})=0$, we have
-(4) $(f_{*}O_{Z}(mE_{Z}))^{\wedge}=O_{\mathbb{P}_{A}}$
-for all sufficiently divisible $m>0$, where $\wedge$ means taking reflexive hull.
-By Gongyo’s abundance theorem and Hacon-Xu [HX13, Theorem 2.12], $(Z,D_{Z})$ has a good minimal model over a nonempty Zariski open subset $U$ of $\mathbb{P}_{A}$.
-We run a $(Z,D_{Z})$-MMP over $\mathbb{P}_{A}$ with scaling of an ample divisor.
-Then this MMP will terminate over $U$ by [HX13, Corollary 2.9].
-Hence we have a partial MMP $(Z,D_{Z})\dashrightarrow(Z^{\prime},D_{Z^{\prime}})$ with $K_{Z^{\prime}}+D_{Z^{\prime}}$ semiample over $U$, where $D_{Z^{\prime}}$ is the pushforward of $D_{Z}$.
-Let $F^{\prime}$ be a general fiber of $f^{\prime}:Z^{\prime}\to\mathbb{P}_{A}$ and let $E_{Z^{\prime}}$ be the pushforward of $E_{Z}$.
-Since $\kappa(K_{F^{\prime}}+D_{Z^{\prime}}|_{F^{\prime}})=0$, we have $E_{Z^{\prime}}|_{F^{\prime}}\sim_{\mathbb{Q}}K_{Z^{\prime}}+D_{Z^{\prime}}|_{F^{\prime}}\sim_{\mathbb{Q}}0$, so $E_{Z^{\prime}}$ is vertical over $\mathbb{P}_{A}$.
-By formula (4), we have $(f^{\prime}_{*}O_{Z^{\prime}}(mE_{Z^{\prime}}))^{\wedge}=(f_{*}O_{Z}(mE_{Z}))^{\wedge}=O_{\mathbb{P}_{A}}$ for all sufficiently divisible $m>0$.
-Hence $E_{Z^{\prime}}$ is very exceptional very $\mathbb{P}_{A}$ by [GL13, Lemma 2.15].
-Then $(Z^{\prime},D_{Z^{\prime}})$, and so $(Z,D_{Z})$, has a good minimal model over $\mathbb{P}_{A}$ by Birkar [Bir12, Theorem 1.8].
+(3) By (2) and the assumption in (3), we known that the vertical part $E_{Z}^{v}$ of $E_{Z}$ is degenerate over $\mathbb{P}_{A}$. Since $\kappa(E_{Z}|_{F})=\kappa(K_{F}+D_{Z}|_{F})=0$, we have
+(3.2.1) $O_{\mathbb{P}_{A}}\subset f_{*}O_{Z}(mE_{Z})\subset(f_{*}O_{Z}(mE_{Z}))^{\wedge}=O_{\mathbb{P}_{A}}$
+for all sufficiently divisible $m>0$, where $\wedge$ means taking reflexive hull. In particular,
+$\bigoplus_{m\geq 0}f_{*}O_{Z}(m(K_{Z}+D_{Z}))=\bigoplus_{m\geq 0}f_{*}O_{Z}(mE_{Z})$
+is a finitely generated $O_{\mathbb{P}_{A}}$-algebra.
+By Gongyo’s abundance theorem and Hacon-Xu [hacon2013existence, Theorem 2.12], $(Z,D_{Z})$ has a good minimal model $(W,D_{W})$ over $\mathbb{P}_{A}$.
+Denote by $g:W\to\mathbb{P}_{A}$ the induced fibration.
+Let $m>0$ be a sufficiently divisible integer.
+Then there is a surjective morphism
+$g^{*}g_{*}O_{W}(m(K_{W}+D_{W}))\to O_{W}(m(K_{W}+D_{W})).$
+By formula (3.2.1), we have
+$g_{*}O_{W}(m(K_{W}+D_{W}))=f_{*}O_{Z}(m(K_{Z}+D_{Z}))=O_{\mathbb{P}_{A}}.$
+Hence $K_{W}+D_{W}\sim_{\mathbb{Q}}0$.
 (4) Write the canonical bundle formula for $g:(W,D_{W})\to\mathbb{P}_{A}$ as $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+B_{\mathbb{P}_{A}}+M_{\mathbb{P}_{A}})$.
-Since $f^{-1}H\subset D_{Z}$, we have $g^{-1}H\subset D_{W}$, so $B_{\mathbb{P}_{A}}\geq H$.
-Since $K_{W}+D_{W}\sim_{\mathbb{Q}}0$ and $K_{\mathbb{P}_{A}}+H\sim 0$, we have $B_{\mathbb{P}_{A}}=H$ and $M_{\mathbb{P}_{A}}\sim_{\mathbb{Q}}0$, so the canonical bundle formula writes
-(5) $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H).$
+Since $f^{-1}H\subset\lfloor D_{Z}\rfloor$, we have $g^{-1}H\subset\lfloor D_{W}\rfloor$.
+Hence $B_{\mathbb{P}_{A}}\geq H$.
+Since $K_{W}+D_{W}\sim_{\mathbb{Q}}0$ and $K_{\mathbb{P}_{A}}+H\sim 0$, we have $B_{\mathbb{P}_{A}}=H$ and $M_{\mathbb{P}_{A}}\sim_{\mathbb{Q}}0$.
+Hence the canonical bundle formula writes
+(3.2.2) $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H).$
 Let $\Sigma$ be a reduced snc divisor on $G$ and let $\overline{\Sigma}$ be its closure in $\mathbb{P}_{A}$.
@@
 By inversion of adjunction (cf.
-[Fil20, Proposition 4.16]), $(W,D_{W}+g^{*}\overline{\Sigma})|_{g^{-1}G}$ is lc.
-Hence the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable by [Kol23, Corollary 4.55].
-In particular, the restriction $g|_{g^{-1}G}$ is equidimensional, and so $W$ has no exceptional divisors over $G$.
-Let $Q\subset D_{W}^{v}$ be a prime divisor.
-If $g(Q)\not\subset H$, then $g(Q)$ is a divisor on $\mathbb{P}_{A}$.
-But by the canonical bundle formula (5), we must have $H\geq g(Q)$, which is a contradiction.
-Hence $D_{W}^{v}=g^{-1}H$.
-(5) By (3), it suffices to show that for any prime divisor $P\subset H$, there is a prime divisor $Q\not\subset E_{Z}$ dominating $P$.
+[Ambro2004Shokurov, Theorem 3.1] or [filipazzi2018generalized, Proposition 4.16]), $(W,D_{W}+g^{*}\overline{\Sigma})|_{g^{-1}G}$ is lc.
+Hence the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable by [Kollar-Families, Corollary 4.55].
+In particular, by [Kollar-Families, Definition-Theorem 4.7], the restriction $g|_{g^{-1}G}$ is flat.
+Let $Q\subset D_{W}^{v}$ be a prime divisor. If $g(Q)\not\subset H$, then $P:=g(Q)$ is a divisor on $\mathbb{P}_{A}$. Since $(W,D_{W}+g^{*}P)$ is lc near the generic point of $P$, we have
+$1\leq a(Q,W,D_{W}+g^{*}P)=-\operatorname{mult}_{Q}(D_{W}+g^{*}P)=-1-\operatorname{mult}_{Q}D_{W}<-1,$
+which is a contradiction. Hence $D_{W}^{v}\subset g^{-1}H\subset\lfloor D_{W}^{v}\rfloor$.
+Since the birational map $Z\dashrightarrow W$ contracts precisely the divisor $E_{Z}\subset\mathrm{Ex}(\pi)$, we have $W\dashrightarrow X$ is a birational contraction and $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$.
+Let $X\xleftarrow{u}V\xrightarrow{v}W$ be a common log resolution of $(X,D_{X})$ and $(W,D_{W})$.
+Then $u^{*}(K_{X}+D_{X})=v^{*}(K_{W}+D_{W})$ by the Negativity Lemma (cf.
+[KM98, 3.39]), i.e., $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ is crepant.
+(5) Case 1.
+$\dim\mathbb{P}_{A}-\dim A=0$.
+By Gongyo’s abundance theorem, $(Z,D_{Z})$ has a good minimal model $(W,D_{W})$.
+Let $Z\xleftarrow{u}V\xrightarrow{v}W$ be a common resolution.
+By [hacon2007shokurov], the rigidity Lemma, and the universal property for Albanese morphisms, we have the following commutative diagram.
+[…]
+where $p_{Z}$, $p_{W}$, $p_{Z}\circ u$, $p_{W}\circ v$ are Albanese morphisms, $A_{Z}\to A$ is the induced morphism by the universal property of $p_{Z}$.
+Hence we may identify the Albanese varieties of $W$ and $Z$, and there is a morphism $W\to A$, compatible with the rational map $Z\dashrightarrow W$.
+Hence $(W,D_{W})$ is actually a good minimal model of $(Z,D_{Z})$ over $A$.
+Case 2.
+$\dim\mathbb{P}_{A}-\dim A>0$.
+By (3),to prove (5), it suffices to show that for any prime divisor $P\subset H$, there is a prime divisor $Q\not\subset E_{Z}$ dominating $P$.
 Assume the contrary, i.e., assume that there is a prime divisor $P\subset H$, and an effective exceptional over $\mathbb{P}_{A}$ divisor $E$ with $f(E)\subset P$ such that
-(6) $f^{-1}P\subset E_{Z}+E.$
-Case 1.
-$\dim\mathbb{P}_{A}-\dim A=1$.
-By [BFPT24, Theorem 1.5], the Albanese morphism $p_{X}:X\to A^{\prime}$ of $X$ is a fibration.
-Since $X$ has rational singularities, we have $\dim A^{\prime}=q(X)=q(Z)=\dim A$.
-Hence, the induced morphism $A\to A^{\prime}$ is an isomorphism.
-We identify $A^{\prime}$ with $A$.
-We have the following commutative diagram.
-[…]
-where $p$ is the natural projection.
-We claim that
-(7) $f^{-1}P\subset E_{Z}.$
-Let $Q\subset E$ be a prime divisor.
-Then $Q\subset f^{-1}H\subset D_{Z}$.
-Since $\dim\mathbb{P}_{A}-\dim A=1$, $f(Q)$ is vertical over $A$, so $f(Q)$ is not an lc center of $(X,D_{X})$ by [BFPT24, Theorem 1.5].
-Hence
-$1<a(Q,X,D_{X})=-\operatorname{mult}_{Q}(D_{Z}-E_{Z})=-1+\operatorname{mult}_{Q}E_{Z}.$
-Hence $Q\subset E_{Z}$ and the claim follows.
-Let $L_{A}$ be an effective and sufficiently ample divisor on $A$. Since $P$ is ample over $A$, we have
-$\dim\mathbb{P}_{A}$ $=\kappa(P+p^{*}L_{A})$
-$=\kappa(f^{*}P+(p\circ f)^{*}L_{A})$
-[…]
-$=\kappa(E_{Z}+(p_{X}\circ\pi)^{*}L_{A})$
-$=\kappa(p_{X}^{*}L_{A})\text{ ($E_{Z}$ is exceptional over $X$)}$
-$=\dim A,$
-which is a contradiction, so the inclusion (6) can not be true.
-Case 2.
-$\dim\mathbb{P}_{A}-\dim A>1$.
-Since $P\subset H=\sum_{i}p_{i}^{*}H_{i}$, we have $P\subset p_{i}^{*}H_{i}$ for some $i$.
-By re-indexing $i$’s, we may assume that $P\subset p_{1}^{*}H_{1}$.
-Let
-$\mathbb{P}^{\prime}_{A}:=\mathbb{P}_{A}(O_{A}\oplus L_{2})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}),$
-let $p_{i}^{\prime}:\mathbb{P}_{A}^{\prime}\to\mathbb{P}_{A}(O_{A}\oplus L_{i})$, $2\leq i\leq d$, $\alpha:\mathbb{P}_{A}\to\mathbb{P}_{A}^{\prime}$ be the natural projections and let $H^{\prime}=\sum_{i=2}^{d}{p^{\prime}}_{i}^{*}H_{i}$.
-Then
-$(\alpha\circ f)^{-1}H^{\prime}\subset f^{-1}H\subset D_{Z}.$
+(3.2.3) $f^{-1}P\subset E_{Z}+E.$
+Let $L_{i}\in\text{Pic}^{0}(A)$, $1\leq i\leq d$, be line bundles such that $G$ is the fiber product over $A$ of the total spaces of $L_{i}$’s with their zero sections removed.
+Then $\mathbb{P}_{A}=\mathbb{P}_{A}(O_{A}\oplus L_{1}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1})$.
+Let $(z_{1},\cdots,z_{d})$ be the inhomogeneous fiber coordinates of $\mathbb{P}_{A}$.
+For each $i$, let $p_{i}:\mathbb{P}_{A}\to\mathbb{P}_{A}(O_{A}\oplus L_{i}^{-1})$ be the projection, and let
+$H_{i}=(z_{i}=0)+(z_{i}=\infty)\subset\mathbb{P}_{A}(O_{A}\oplus L_{i}^{-1})$
+be the effective and ample over $A$ divisor. Then $H=\sum_{i=1}^{d}p_{i}^{*}H_{i}$.
+Since $P\subset H=\sum_{i=1}^{d}p_{i}^{*}H_{i}$, we have $P\subset p_{i}^{*}H_{i}$ for some $i$. By re-indexing $i$’s, we may assume that $P\subset p_{1}^{*}H_{1}$. Let
+$\mathbb{P}^{\prime}_{A}=\mathbb{P}_{A}(O_{A}\oplus L_{2}^{-1})\times_{A}\cdots\times_{A}\mathbb{P}_{A}(O_{A}\oplus L_{d}^{-1}),$
+and let $p_{i}^{\prime}:\mathbb{P}_{A}^{\prime}\to\mathbb{P}_{A}(O_{A}\oplus L_{i}^{-1})$, $2\leq i\leq d$, $\alpha:\mathbb{P}_{A}\to\mathbb{P}_{A}^{\prime}$, $p^{\prime}:\mathbb{P}_{A}^{\prime}\to A$ be the natural projections.
+Let $H^{\prime}=\sum_{i=2}^{d}{p^{\prime}}_{i}^{*}H_{i}$.
+Then $\mathbb{P}_{A}^{\prime}\setminus H^{\prime}$ is a quasi-abelian variety whose canonical compactification is just $(\mathbb{P}_{A}^{\prime},H^{\prime})$, and $\alpha\circ f:Z\to\mathbb{P}_{A}^{\prime}$ is a fibration with
+$(\alpha\circ f)^{-1}H^{\prime}\subset f^{-1}H\subset\lfloor D_{Z}\rfloor.$
 By induction on $(\dim\mathbb{P}_{A}-\dim A)$, $(Z,D_{Z})$ has a good minimal model $(V,D_{V})$ over $\mathbb{P}_{A}^{\prime}$ with $K_{V}+D_{V}\sim_{\mathbb{Q}}0$.
@@
 […]
-where $p^{\prime}$ is the natural projection and $p_{V}$ is the Albanese morphism of $V$ (by the same reasoning as in Case 1, we may identify the Albanese varieties of $Z$ and $V$).
 We claim that
-(8) $f^{-1}P\subset E_{Z}+(\alpha\circ f)^{-1}H^{\prime}.$
+(3.2.4) $f^{-1}P\subset E_{Z}+(\alpha\circ f)^{-1}H^{\prime}.$
 Let $Q\subset E$ be a prime divisor which is not contained in $E_{Z}$.
-Then $Q\subset f^{-1}H\subset D_{Z}$.
-Since $Z\dashrightarrow V$ contracts precisely the divisor $E_{Z}$, the strict transform $Q_{V}$ of $Q$ on $V$ is a divisor contained in $D_{V}$.
+Since $f(E)\subset P\subset H$, we have $Q\subset f^{-1}H\subset\lfloor D_{Z}\rfloor$.
+Since the MMP $Z\dashrightarrow V$ contracts precisely the divisor $E_{Z}$, the strict transform $Q_{V}$ of $Q$ on $V$ is a divisor contained in $\lfloor D_{V}\rfloor$.
 Since $\dim\mathbb{P}_{A}-\dim\mathbb{P}_{A}^{\prime}=1$, and $Q$ is exceptional over $\mathbb{P}_{A}$, we have $Q$ is vertical over $\mathbb{P}_{A}^{\prime}$.
 Hence $Q_{V}$ is vertical over $\mathbb{P}_{A}^{\prime}$.
-By (4), we have $g(Q_{V})\subset H^{\prime}$.
+By (4), we have $Q_{V}\subset g^{-1}H^{\prime}$.
 Hence $Q\subset(\alpha\circ f)^{-1}H^{\prime}$ and the claim follows.
@@
 $=\dim\mathbb{P}_{A}^{\prime},$
-which is a contradiction, so the inclusion (6) can not be true. ∎
-Lemma 16.
+which is a contradiction. Hence the inclusion (3.2.3) can not be true. ∎
+Lemma 3.3 (cf. [Deng2024HyperFund, Proposition 1.3]).
 Let $G$ be a quasi-abelian variety and let $P\subset G$ be a prime divisor, then $\overline{\kappa}(G\setminus P)>0$.
 Proof.
-The idea is from [Fuj24, Proof of Theorem 1.3, Step 2]. Let $A$ be the abelian part of $G$ and let $p:G\to A$ be the natural projection.
+The idea is from [fujino2024quasi, Proof of Theorem 1.3, Step 2]. Let $A$ be the compact part of $G$ and let $p:G\to A$ be the natural projection.
 Case 1.
@@
 Then $P=p^{*}Q$ for some prime divisor $Q$ on $A$.
-Let $\mathbb{P}_{A}\to A$ be a $(\mathbb{P}_{1})^{d}$-bundle over $A$ such that $\mathbb{P}_{A}$ is a compactification of $G$, and $H=\mathbb{P}_{A}\setminus G$ is a snc divisor with $K_{\mathbb{P}_{A}}+H\sim 0$.
+Let $(\mathbb{P}_{A},H)$ be the canonical compactification of $G$.
 Let $\overline{P}$ be the closure of $P$ in $\mathbb{P}_{A}$.
 Since any lc center of $(\mathbb{P}_{A},H)$ is horizontal over $A$, $\overline{P}$ contains no lc centers of $(\mathbb{P}_{A},H)$.
-Then $\overline{\kappa}(G\setminus P)=\kappa(K_{\mathbb{P}_{A}}+H+\overline{P})=\kappa(\overline{P})=\kappa(Q)>0$ by Lemma 17.
+Then $\overline{\kappa}(G\setminus P)=\kappa(K_{\mathbb{P}_{A}}+H+\overline{P})=\kappa(\overline{P})=\kappa(Q)>0$ by Lemma 3.4.
 Case 2.
@@
 Then $F$ is a curve with $\overline{\kappa}(F)>0$.
-Hence $\overline{\kappa}(G\setminus P)\geq\overline{\kappa}(F)+\overline{\kappa}(G^{\prime})=\overline{\kappa}(F)>0$ by Kawamata [Kaw78, Theorem 1] (see also Fujino [Fuj20, Chapter 5]).∎
-Lemma 17.
+Hence $\overline{\kappa}(G\setminus P)\geq\overline{\kappa}(F)+\overline{\kappa}(G^{\prime})=\overline{\kappa}(F)>0$ by Kawamata [kawamata1977addition, Theorem 1] (see also Fujino [fujino2020iitaka, Chapter 5]).∎
+Lemma 3.4.
 Let $(Y,D_{Y})$ be a projective log smooth pair with $D_{Y}$ reduced.
@@
 $=\kappa(K_{Y}+D_{Y}+\Sigma)\text{ (by $\kappa(K_{Y}+D_{Y})\geq 0$)}.$
-Lemma 18.
+Lemma 3.5.
 Let $\alpha:Z\to V$ and $\beta:W\to V$ be fibrations between projective normal varieties, and let $Z\dashrightarrow W$ be a birational contraction over $V$.
@@
 $\kappa(\alpha^{*}L+E)=\kappa(u^{*}\alpha^{*}L+u^{*}E)=\kappa(v^{*}\beta^{*}L+u^{*}E)=\kappa(\beta^{*}L)=\kappa(L).$
-Theorem 2 is a special case of the following theorem.
-Theorem 19.
-Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$.
-Let $\alpha:X\setminus D\to G$ be the quasi-Albanese morphism.
-Then the conclusion in Theorem 2 holds.
+Theorem 3.6 (A stronger dlt version of Theorem 1.4).
+Let $(X,D_{X})$ be a projective $\mathbb{Q}$-factorial dlt pair with $K_{X}+D_{X}\sim_{\mathbb{Q}}0$.
+Let $\Delta_{X}\leq\lfloor D_{X}\rfloor$ be a reduced divisor.
+Let $\alpha:X\setminus\Delta_{X}\to G$ be the quasi-Albanese morphism.
+Let $A$ be the compact part of $G$.
+Let $(\mathbb{P}_{A},H)$ be the canonical compactification of $G$.
+Then there are
+  • a projective $\mathbb{Q}$-factorial dlt pair $(W,D_{W})$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$ and
+  • a crepant birational contraction $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ with $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$,
+such that the following holds.
+  1. (1)
+    The rational map $g:W\dashrightarrow\mathbb{P}_{A}$ is a fibration,
+  2. (2)
+    The canonical bundle formula for $(W,D_{W})\to\mathbb{P}_{A}$ writes $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H)$, and the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable.
+  3. (3)
+    Let $\Delta_{W}=\mu^{-1}_{*}\Delta_{X}+\operatorname{Ex}(\mu)$.
+Then $\Delta^{v}_{W}=D_{W}^{v}=\lfloor D_{W}^{v}\rfloor=g^{-1}H$ and the morphism $W\setminus\Delta_{W}\to G$ is the quasi-Albanese morphism of $W\setminus\Delta_{W}$.
 Proof.
-Keep the notations introduced in the proof of Theorem 3.
+Let $\pi:Z\to X$ be a log resolution such that the rational map $f:Z\dashrightarrow\mathbb{P}_{A}$ is a morphism.
+By Theorem 3.1(2), $f$ is a fibration.
+Let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$ and let $\Delta_{Z}=\pi^{-1}_{*}\Delta_{X}+\operatorname{Ex}(\pi)$.
+Then $K_{Z}+D_{Z}=\pi^{*}(K_{X}+D_{X})+E_{Z}$ for some effective and exceptional divisor $E_{Z}$ and $f^{-1}H\subset\Delta_{Z}\subset\lfloor D_{Z}\rfloor$.
+By Theorem 3.2, there is a good minimal model $(W,D_{W})$ of $(Z,D_{Z})$ over $\mathbb{P}_{A}$ satisfying (1) and (2) and such that the birational map $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ is a crepant birational contraction with $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$.
+By $\lfloor D_{W}^{v}\rfloor=g^{-1}H\subset\Delta_{W}^{v}\subset\lfloor D_{W}^{v}\rfloor$, we have $\Delta_{W}^{v}=\lfloor D_{W}^{v}\rfloor$.
+Let $W\setminus\Delta_{W}\to G^{\prime}$ be the quasi-Albanese morphism.
+Since $W\setminus\Delta_{W}\to G$ is a fibration, the induced morphism $G^{\prime}\to G$ is a fibration.
+By Lemmas 2.3 and 2.7, we have $\dim G^{\prime}=q(W,\Delta_{W})=q(X,\Delta_{X})=\dim G$.
+Hence the fibration $G^{\prime}\to G$ is an isomorphism, i.e., $W\setminus\Delta_{W}\to G$ is the quasi-Albanese morphism.
+∎
+Proof of Theorem 1.4.
+Let $\pi:Z\to X$ be a log resolution such that $\pi^{-1}\operatorname{nklt}$ is a divisor and the rational map $f:Z\dashrightarrow\mathbb{P}_{A}$ is a morphism.
+By Theorem 3.1(3), $f$ is a fibration.
+Let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$.
+Then $K_{Z}+D_{Z}=\pi^{*}(K_{X}+D_{X})+E_{Z}$ for some effective and exceptional divisor $E_{Z}$ and $f^{-1}H\subset\pi^{-1}\operatorname{nklt}\subset\lfloor D_{Z}\rfloor$.
+By Theorem 3.2, there is a good minimal model $(W,D_{W})$ of $(Z,D_{Z})$ over $\mathbb{P}_{A}$ satisfying (1) and (2) and such that the birational map $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ is a crepant birational contraction with $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$.
+By the same reasoning as in the previous Theorem 3.6, we have $W\setminus\lfloor D_{W}\rfloor\to G$ is the quasi-Albanese morphism. ∎
+Proof of Theorem 1.3.
+Keep the notations introduced in the proof of Theorem 1.4.
 (1) Let $P\subset G$ be a prime divisor and let $\overline{P}$ be its closure in $\mathbb{P}_{A}$.
 Then $\overline{P}\not\subset H$.
-Since $D_{W}^{v}=g^{-1}H$, there is a prime divisor $\overline{Q}_{W}\subset W$ such that $\overline{Q}_{W}\not\subset D_{W}$ and $\overline{Q}_{W}$ dominates $\overline{P}$.
-Let $\overline{Q}$ be the strict transform of $\overline{Q}_{W}$ on $X$.
-Then $\overline{Q}\not\subset D_{X}$ and $\overline{Q}$ dominates $P$.
-(2) Let $Q\subset X\setminus D_{X}$ be a prime divisor and let $\overline{Q}$ be its closure in $X$.
-Then $\overline{Q}\not\subset D_{X}$ and its strict transform $\overline{Q}_{W}$ on $W$ has $\overline{Q}_{W}\not\subset D_{W}$, so $\overline{Q}_{W}\not\subset g^{-1}H$.
-Since $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable and in particular equidimensional, we have $\operatorname{codim}\alpha(Q)=\operatorname{codim}g(\overline{Q}_{W})=1$.
-(3) Let $P\subset G$ be a prime divisor and let $\overline{P}$ be its closure in $\mathbb{P}_{A}$.
-Since $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, we have $(W,D_{W}+g^{*}\overline{P})$ is lc near the generic point of $\overline{P}$, so $(Z,D_{Z}-E_{Z}+f^{*}\overline{P})|_{f^{-1}G}$ is sub-lc near the generic point of $P$.
+Since $\lfloor D_{W}^{v}\rfloor=g^{-1}H$, there is a prime divisor $\overline{Q}_{W}\subset W$ such that $\overline{Q}_{W}\not\subset\lfloor D_{W}\rfloor$ and $\overline{Q}_{W}$ dominates $\overline{P}$.
+Since $\operatorname{Ex}(\mu)\leq\lfloor D_{W}\rfloor$, $\overline{Q}:=\mu_{*}\overline{Q}_{W}$ is a prime divisor on $X$ such that $\overline{Q}\not\subset\lfloor D_{X}\rfloor$ (in particular $\overline{Q}\not\subset\operatorname{nklt}(X,D_{X})$) and $\overline{Q}$ dominates $P$.
+(2) Let $\Sigma\subset G$ be a reduced snc divisor and let $\overline{\Sigma}$ be its closure in $\mathbb{P}_{A}$.
+Since $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, we have $(W,D_{W}+g^{*}\overline{\Sigma})|_{g^{-1}G}$ is lc.
+Hence $(Z,D_{Z}-E_{Z}+f^{*}\overline{\Sigma})|_{f^{-1}G}$ is sub-lc.
 By
-$(K_{Z}+D_{Z}-E_{Z}+f^{*}\overline{P})|_{Z\setminus\pi^{-1}D_{X}}=(\pi|_{Z\setminus\pi^{-1}D_{X}})^{*}(K_{X\setminus D_{X}}+\alpha^{*}P)$
-and $Z\setminus\pi^{-1}D_{X}\subset Z\setminus f^{-1}H=f^{-1}{G}$, we have $(X\setminus D_{X},\alpha^{*}P)$ is lc near the generic point of $P$. ∎
-4. A remark on orbifold fundamental groups
-This section is inspired by the work [GLM23].
-Let $(X,D_{X}=\sum_{i}(1-\frac{1}{m_{i}})D_{i})$ be an lc pair with standard coefficients, where each $D_{i}$ is a prime divisor and $m_{i}\in\mathbb{Z}_{>0}\cup\infty$.
-For each $i$, let $\gamma_{i}$ be a small loop around $D_{i}$.
-The orbifold fundamental group $\pi_{1}^{\operatorname{orb}}(X,D_{X})$ is defined as the quotient of $\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})$ by the normal subgroup generated by $\gamma_{i}^{m_{i}}$’s.
-Note that if $D_{X}$ is reduced, then $\pi_{1}^{\operatorname{orb}}(X,D_{X})=\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})$.
-Conjecture 20.
-Let $(X,D_{X})$ be an lc pair with standard coefficients. If $K_{X}+D_{X}\sim_{\mathbb{Q}}0$, then $\pi_{1}^{\operatorname{orb}}(X,D)$ is virtually nilpotent.
-Using the index-one cover and the Galois correspondence for orbifold fundamental groups (cf.
-[GLM23, Proposition A.11]), Conjecture 20 is reduced to the case $D_{X}$ is reduced.
-By the following Lemma, Conjecture 20 is further reduced to the $\mathbb{Q}$-factorial dlt case.
-Lemma 21.
-Let $(X,D_{X})$ and $(Y,D_{Y})$ be two projective lc pairs with $D_{X}$ and $D_{Y}$ reduced.
-If there is a birational contraction $p:X\dashrightarrow Y$ such that $D_{X}=p^{-1}_{*}D_{Y}+\operatorname{Ex}(p)$, then there is an isomorphism $\pi_{1}^{\operatorname{orb}}(X,D_{X})\cong\pi_{1}^{\operatorname{orb}}(Y,D_{Y})$.
+$(K_{Z}+D_{Z}-E_{Z}+f^{*}\overline{\Sigma})|_{Z\setminus\pi^{-1}\operatorname{nklt}}=(\pi|_{Z\setminus\pi^{-1}\operatorname{nklt}})^{*}(K_{X\setminus\operatorname{nklt}}+D_{X}|_{X\setminus\operatorname{nklt}}+\alpha^{*}\Sigma)$
+and $Z\setminus\pi^{-1}\operatorname{nklt}\subset Z\setminus f^{-1}H=f^{-1}{G}$, we have $(X\setminus\operatorname{nklt}+D_{X}|_{X\setminus\operatorname{nklt}},\alpha^{*}\Sigma)$ is lc.
+Hence $\alpha:(X\setminus\operatorname{nklt},D_{X}|_{X\setminus\operatorname{nklt}})\to G$ is locally stable by [Kollar-Families, Corollary 4.55].
+It is flat with slc fibers by [Kollar-Families, Definition-Theorem 4.7]
+3.3. Proofs of Lemma 1.5 and Theorem 1.6
+Lemma 3.7.
+Let $(X,D_{X})$ be a projective log smooth pair with $D_{X}$ reduced. Write $D_{X}=\sum_{i=1}^{m}D_{i}$ as a sum of prime divisors. Then
+$q(X,D_{X})-q(X)=\{(a_{1},\cdots,a_{m})\in\mathbb{C}^{m}\mid\sum_{i}a_{i}c_{1}(O_{X}(D_{i}))=0\text{ in }H^{1}(\Omega_{X}^{1})\},\text{ and}$
+[…]
 Proof.
-Let $Z\subset X$ and $Z^{\prime}\subset Y$ be closed subsets such that $X\setminus Z\cong Y\setminus Z^{\prime}$.
-Hence $X^{\operatorname{reg}}\setminus Z\cong Y^{\operatorname{reg}}\setminus Z^{\prime}$ and $X^{\operatorname{reg}}\setminus D_{X}\cup Z\cong Y^{\operatorname{reg}}\setminus D_{Y}\cup Z^{\prime}$.
-Since $p$ is a birational contraction, we have $\operatorname{codim}Z^{\prime}\geq 2$ and $D_{X}\cup Z=D_{X}\cup Z^{\prime\prime}$ for some $\operatorname{codim}\geq 2$ closed subset $Z^{\prime\prime}$.
-Hence $\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})\cong\pi_{1}(X^{\operatorname{reg}}\setminus D_{X}\cup Z^{\prime\prime})\cong\pi_{1}(X^{\operatorname{reg}}\setminus D_{X}\cup Z)\cong\pi_{1}(Y^{\operatorname{reg}}\setminus D_{Y}\cup Z^{\prime})\cong\pi_{1}(Y^{\operatorname{reg}}\setminus D_{Y})$.
+Consider the exact sequence
+$0\to\Omega_{X}^{1}\to\Omega_{X}^{1}(\log D_{X})\to\bigoplus_{i=1}^{m}O_{D_{i}}\to 0,$
+and the associated long exact sequence:
+$0\to H^{0}(\Omega_{X}^{1})\to H^{0}(\Omega_{X}^{1}(\log D_{X}))\to\bigoplus_{i=1}^{m}H^{0}(O_{D_{i}})\xrightarrow{\delta}H^{1}(\Omega_{X}^{1}).$
+Since $\delta(1|_{D_{i}})=c_{1}(O_{X}(D_{i}))\in H^{1}(\Omega_{X}^{1})$ (cf.
+[HartshorneGTM52, Chapter III Exercise 7.4]), where $1|_{D_{i}}$ is the constant function on $D_{i}$ with value $1$, the Lemma follows.
 ∎
-Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$.
-By a theorem due to Arapura and Nori [AN99, Corollary 4.10], to show that $\pi_{1}^{\text{orb}}(X,D_{X})=\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})$ is virtually nilpotent, it suffices to show that it is polycyclic.
-Recall that a group $G$ is called polycyclic if there is a series of subgroups $1=G_{n}\triangleleft G_{n-1}\triangleleft\cdots\triangleleft G_{0}=G$ such that each $G_{i}/G_{i+1}$ is cyclic.
-The advantage of the class of polycyclic groups is that they are closed under passing to subgroups, quotients, and extensions (cf.
-[KM79, 4.4.3.
-Exercise]), which is crucial to inductive arguments.
-Assume that $q(X,D_{X})>0$, and let $g:(W,D_{W})\to\mathbb{P}_{A}$ be the fibration given by Theorem 3.
-By Lemma 21, $(X,D_{X})$ and $(W,D_{W})$ have the same orbifold fundamental group.
-Let $W^{0}=g^{-1}G$ and let $F$ be a general fiber of $g$.
-Then $(F,{D_{W}}|_{F})$ is lc and $K_{F}+D_{W}|_{F}\sim_{\mathbb{Q}}0$.
-By the following exact sequence (cf.
-[GLM23, Lemma 2.18])
-$\pi_{1}^{\operatorname{orb}}(F,D_{W}|_{F})\to\pi_{1}^{\operatorname{orb}}(W^{0},{D_{W}}|_{W^{0}})\to\pi_{1}(G)\to 1,$
-we have $\pi_{1}^{\operatorname{orb}}(X,D_{X})\cong\pi_{1}^{\operatorname{orb}}(W,D_{W})\cong\pi_{1}^{\operatorname{orb}}(W^{0},{D_{W}}|_{W^{0}})$ is polycyclic if $\pi_{1}^{\operatorname{orb}}(F,D_{W}|_{F})$ is polycyclic.
-Hence, we have shown the following by induction on dimension.
-Proposition 22.
-To prove Conjecture 20, it suffices to prove $\mathbb{Q}$-factorial dlt pairs $(X,D_{X})$ with $D_{X}$ reduced and $q(X,D_{X})=0$ have polycyclic orbifold fundamental groups.
+Lemma 3.8 (cf. [enwright2024log, Lemma 2.41]).
+Let $(X,D_{X})$ be a projective lc pair. Let $\pi:Z\to X$ be a log resolution and let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Then $c(Z,D_{Z})=c(X,D_{X})$.
+Proof.
+Write $\operatorname{Ex}(\pi)=\sum_{i=1}^{m}E_{i}$ as a sum of prime divisors.
+Since $|\lfloor D_{Z}\rfloor|=|\lfloor D_{X}\rfloor|+m$, to prove the Lemma, it suffices to show that $\operatorname{rank}\langle\lfloor D_{Z}\rfloor\rangle=\operatorname{rank}\langle\lfloor D_{X}\rfloor\rangle+m$.
+Since $\pi_{*}:\langle\lfloor D_{Z}\rfloor\rangle\otimes\mathbb{Q}\to\langle\lfloor D_{X}\rfloor\rangle\otimes\mathbb{Q}$ is surjective and $E_{i}$’s form a basis of its kernel (by the negativity Lemma [KM98, 3.39]), the Lemma follows.
+∎
+Proof of Lemma 1.5.
+Let $\pi:Z\to X$ be a log resolution such that $\pi^{-1}\operatorname{nklt}(X,D_{X})$ is a divisor.
+Let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$.
+By the previous two Lemmas, we have $c(X,D_{X})=c(Z,D_{Z})=\dim Z-(q(Z,D_{Z})-q(Z))$.
+By Lemma 2.7, $X\setminus\operatorname{nklt}(X,D_{X})$ and $Z\setminus\lfloor D_{Z}\rfloor$ have the isomorphic quasi-Albanese variety $G$.
+By Iitaka’s construction, we have $\dim G=q(Z,D_{Z})$ and $\dim A=q(Z)$.
+Hence $c(X,D_{X})=\dim X-(\dim G-\dim A)$.
+∎
+Proof of Theorem 1.6.
+Let $\Delta_{X}\leq\lfloor D_{X}\rfloor$ be a reduced divisor.
+Let $\pi:Z\to X$ be a Hacon’s $\mathbb{Q}$-factorial dlt modification [Hacon2014ACC, Proposition 3.3.1].
+Let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$ and let $\Delta_{Z}=\pi^{-1}_{*}\Delta_{X}+\operatorname{Ex}(\pi)$.
+Then $K_{Z}+D_{Z}=\pi^{*}(K_{X}+D_{X})$, $\kappa(K_{Z}+D_{Z})=0$ and $\Delta_{Z}\leq\lfloor D_{Z}\rfloor$.
+By Lemma 3.8, we have $c(X,\Delta_{X})=c(Z,\Delta_{Z})$.
+By Proposition 1.5, we have $c(Z,\Delta_{Z})=\dim Z-(\dim G-\dim A)$, where $G$ and $A$ are the quasi-Albanese variety and the Albanese variety of $Z\setminus\Delta_{Z}$ and $Z$ respectively.
+By 3.1(2), we have $\dim Z\geq\dim G$.
+Hence $c(X,\Delta_{X})\geq 0$.
+Assume that $c(X,\Delta_{X})=0$.
+Then $\dim Z=\dim G$ and $\dim A=0$.
+Hence $G\cong(\mathbb{C}^{*})^{n}$.
+The quasi-Albanese morphism $Z\setminus\Delta_{Z}\to(\mathbb{C}^{*})^{n}$ is birational, so $X$ is rational.
+Assume further that $K_{X}+D_{X}\sim_{\mathbb{Q}}0$.
+Then $K_{Z}+D_{Z}\sim_{\mathbb{Q}}0$.
+Set $((\mathbb{P}^{1})^{n},H):=(\mathbb{P}^{1}_{z_{1}}\times\cdots\times\mathbb{P}^{1}_{z_{n}},\sum_{i}((z_{i}=0)+(z_{i}=\infty)))$.
+By Theorem 3.6, there are
+  • a $\mathbb{Q}$-factorial dlt pair $(W,D_{W})$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$,
+  • a crepant birational contraction $\mu:(W,D_{W})\dashrightarrow(Z,D_{Z})$ with $D_{W}=\mu^{-1}_{*}D_{Z}+\operatorname{Ex}(\mu)$, and
+  • a birational morphism $g:(W,D_{W})\to((\mathbb{P}^{1})^{n},H)$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{(\mathbb{P}^{1})^{n}}+H)$, and $\Delta_{W}=\lfloor D_{W}\rfloor=D_{W}=g^{-1}H$, where $\Delta_{W}=\mu^{-1}_{*}\Delta_{Z}+\operatorname{Ex}(\mu)$.
+Hence we have $\Delta_{Z}=\lfloor D_{Z}\rfloor=D_{Z}$, $\Delta_{X}=\lfloor D_{X}\rfloor=D_{X}$ and the four pairs $(X,D_{X})$, $(Z,D_{Z})$, $(W,D_{W})$ $((\mathbb{P}^{1})^{n},H)$ are crepant birational to each other.
+Let $P\subset W$ be a prime divisor which is exceptional over $(\mathbb{P}^{1})^{n}$, then $P\subset g^{-1}H=D_{W}$.
+By $1=a(P,W,D_{W})=a(P,(\mathbb{P}^{1})^{n},H)$, we have $P$ is toric.
+By [Brown2018GeomToric, Lemma 2.3.2], we have $(W,D_{W})$, $(Z,D_{Z})$ and $(X,D_{X})$ are toric.
+By taking a log resolution of $(Z,D_{Z})$ and Lemma 2.7, we known that $Z\setminus\lfloor D_{Z}\rfloor$ and $X\setminus\operatorname{nklt}(X,D_{X})$ have the same quasi-Albanese variety $(\mathbb{C}^{*})^{n}$.
+By Theorem 1.3, we have the quasi-Albanese morphism $X\setminus\operatorname{nklt}(X,D_{X})\to(\mathbb{C}^{*})^{n}$ is birational and isomorphic in codimension one.
+∎